### v2

# Preliminary Ranking of WMT25 General Machine Translation Systems

Preliminary Ranking of WMT25 General Machine Translation Systems
[…]
Preliminary Ranking of WMT25 General Machine Translation Systems
Tom Kocmi          Eleftherios Avramidis          Rachel Bawden          Ondřej Bojar          Konstantin Dranch
Anton Dvorkovich          Sergey Dukanov          Natalia Fedorova          Mark Fishel          Markus Freitag
Thamme Gowda          Roman Grundkiewicz          Barry Haddow          Marzena Karpinska
Philipp Koehn          Howard Lakougna          Jessica Lundin          Kenton Murray          Masaaki Nagata
Stefano Perrella          Lorenzo Proietti          Martin Popel          Maja Popović          Parker Riley
Mariya Shmatova          Steinþór Steingrímsson          Lisa Yankovskaya          Vilém Zouhar
Introduction
We present the preliminary rankings of machine translation (MT) systems submitted to the WMT25 General Machine Translation Shared Task,111www2.statmt.org/wmt25/translation-task.html as determined by automatic evaluation metrics. Because these rankings are derived from automatic evaluation, they may exhibit a bias toward systems that employ re-ranking techniques, such as Quality Estimation or Minimum Bayes Risk decoding. The official WMT25 ranking will be based on human evaluation, which is more reliable and will supersede these results. The official WMT25 ranking will be based on human evaluation, which is more reliable and will supersede these results. The purpose of releasing these findings now is to assist task participants with their system description papers; not to provide final findings.
Types of Systems
We distinguish two types of MT systems participating in the shared task:
  • Constrained systems: must use only publicly available training data and models, be limited to a maximum of 20B parameters, and have their model weights released under an open license.
  • Unconstrained systems: (marked with gray) are all other systems. They have no restrictions on training data or model size, and there is no requirement to publish the model weights. This category also includes systems for which training information is not public.
Model # Params Open?
Constrained systems
[…]
Unconstrained systems
[…]
Evaluated Systems
Models.
Our evaluation includes systems submitted by participants, as well as open-weight and proprietary models. We selected the largest or best-performing version of each model where applicable. All constrained and unconstrained systems are listed in Table 1. Full details for all systems will be available in the upcoming WMT25 finding paper.
Prompts.
We prompt all language models using a zero-shot, instruction-following approach, with the specific instructions provided as part of the blind test set. Each model was first tasked with translating the entire document. If this initial attempt failed (e.g., due to producing an incorrect paragraph count or exceeding the token limit), we implemented a fallback strategy of translating the document paragraph by paragraph. This generic setup may disadvantage systems tuned for specific MT instructions, such as TowerLLM or EuroLLM; these are marked with [M].
[…]
Evaluation Data
Languages.
The evaluation covers 32 language pairs, with each test set containing approximately 37K words [^2] As per whitespace split. organized into documents. Half of these language pairs are designated for human evaluation, while the other half belong to a multilingual subtrack evaluated solely with automatic metrics. Most pairs are in the English-to-X direction with the exception of Czech-Ukrainian, Czech-German, and Japanese-Chinese.
Domains.
The test sets combine material from four distinct domains, though language pairs with a non-English source may differ slightly in their domain distribution and dataset size. The domains include:
  • News commentary
  • Social (texts from social networks, collected with screenshots)
  • Speech (transcripts of speeches obtained automatically)
  • Literary (two documents of roughly 5,000 words each)
Data preprocessing.
Each document is divided into segments of approximately 100 words. These segments typically correspond to natural paragraphs; however, some paragraphs were split (or merged) to adhere to this length constraint. Sentence splitting was not performed on the source texts, so segments often contain multiple sentences. For the multimodal context, participants were allowed to (optionally) use accompanying image and video modalities when available, and human annotators who provided reference translations also had access to this context.
All data for this task, including references, system outputs, automatic scores, and the LaTeX source of this document, are publicly available at our repository: github.com/wmt-conference/wmt25-general-mt.
Automatic Ranking
This section details our automatic ranking method, which we refer to as AutoRank. Both the set of automatic metrics and the aggregation procedure have been slightly updated since last year’s shared task.
Metrics.
For most language pairs, [^3] See “Low-resource exception” below. the AutoRank is a combination of three distinct families of evaluation methods:
  • LLM-as-a-Judge (reference-less). We use GEMBA-ESA (Kocmi and Federmann, 2023) with two independent judges: GPT-4.1 (OpenAI, 2025) and Command A (Cohere Team, 2025), both in a reference-less setting.
  • Trained reference-based metrics. Two supervised metrics trained to approximate human quality judgments with references: MetricX-24-Hybrid-XL (Juraska et al., 2024) and XCOMET-XL (Guerreiro et al., 2024).
  • Trained Quality Estimation (QE). The reference-less QE metric CometKiwi-XL (Rei et al., 2023), which is also trained to mimic human judgments.
This combination of reference-based and reference-less (or QE) methods is designed to balance their complementary failure modes. Reference-based metrics typically achieve a higher correlation with human judgments when high-quality references are available, while reference-less methods reduce susceptibility to reference bias when references are suboptimal (Freitag et al., 2023). We also account for known issues with specific metrics. To mitigate a common QE pitfall, i.e., being fooled by fluent output in the wrong language, the GEMBA-ESA prompt explicitly specifies the target language. However, while GEMBA-ESA is intended to reduce bias toward systems that use re-ranking, we note that some participants incorporated it directly as a reward model.
System-level scores.
The system-level score for each language pair is the average of its paragraph-level (segment-level) scores from each metric across the testset. We make one exception for language pairs without human references by excluding CometKiwi-XL from the AutoRank computation. This avoids redundancy, as the other hybrid metrics (MetricX-24-Hybrid-XL and XCOMET-XL) can also run in a reference-less (QE) mode to provide the necessary QE signal.
Low-resource exception.
For the two lowest-resource languages in the testset, i.e., Bhojpuri and Maasai, we rely solely on chrF++ (Popović, 2017), computed with sacrebleu (Post, 2018). This approach was chosen because the reliability of our main metrics is unestablished for these languages (Falcão et al., 2024; Singh et al., 2024; Wang et al., 2024; Sindhujan et al., 2025), whereas high-quality human references required for chrF++ were available.
From system-level scores to AutoRank.
To combine the metrics into a single score, we first normalize them using median-interpercentile scaling to address differences in scale and reduce the influence of low-performing outliers. We then compute the average using equal weights. Finally, we linearly rescale the results to the range from 1 to $N$ systems. A detailed description is provided below:
Let $S$ be the set of submitted systems for a given language pair, $|S|=N$, and let $M$ be the set of automatic metrics used for that language pair (for Bhojpuri and Maasai, $|M|=1$). For each metric $m\in M$ and system $s\in S$, we compute a system-level score $x^{(m)}_{s}$ as the average of that metric over all available test segments. To combine scores across metrics, we first map them to a common scale; however, classical min–max normalization is highly sensitive to outliers. To downweight extremes without discarding any system, we apply a median–interpercentile scaling to each metric $m$:
$\tilde{x}^{(m)}$ $=\operatorname*{median}\,\left\{x^{(m)}_{s}\mid s\in S\right\},$ (1a)
$D^{(m)}$ $=\max\,\left(\varepsilon,\,Q^{(m)}_{100}-Q^{(m)}_{25}\right)$, (1b)
$z^{(m)}_{s}$ $=\frac{x^{(m)}_{s}-\tilde{x}^{(m)}}{D^{(m)}}.$ (1c)
Where $\varepsilon>0$ and $Q^{(m)}_{p}$ denotes the $p$-th percentile of $\{x^{(m)}_{s}:s\in S\}$. Importantly, Eq. (1) is continuous and monotonic: it keeps all systems and preserves their order within each metric. Then, for each system, we average the robust-scaled values across metrics:
$\bar{z}_{s}\;=\;\frac{1}{|M|}\sum_{m\in M}z^{(m)}_{s}$. (2)
Averaging after robust scaling yields a single comparable score that preserves the magnitude of performance differences between systems (in standardized units) while preventing any single metric’s outliers from dominating. Finally, for readability and to follow the WMT convention from last year (lower is better in AutoRank, i.e., $1$ is best and $N$ worst), we apply a final linear mapping to the set $\{\bar{z}_{s}\}_{s\in S}$. Specifically, within $\{\bar{z}_{s}\}_{s\in S}$ the system with the highest average score is assigned $1$, the system with the lowest average score is assigned $N$, and all remaining systems are placed linearly between these two endpoints. This remapping is applied only once—after the cross-metric aggregation—so it preserves the ordering and relative spacing between systems while retaining the outlier mitigation provided by the robust scaling. We refer to the resulting value as AutoRank in the various tables.
Human Evaluation
This year’s shared task saw a record number of participants, with submissions from 36 unique teams. [^4] A total of 43 teams initially registered, but 7 later withdrew or were disqualified. Due to budget constraints, we could not include all submissions in the human evaluation. Therefore, we selected a subset of systems for manual evaluation using the Error Span Annotation (ESA) protocol (Kocmi et al., 2024). This subset typically consists of 18 systems per language pair, although this number is higher for some. For any system not selected for human evaluation, the automatic metric ranking (AutoRank) serves as the official final result.
The selection process for human evaluation prioritized constrained systems over unconstrained ones, following a two-step procedure:
  1. 1.
    First, the top-8 performing constrained systems were selected, based on their automatic scores.
  2. 2.
    Second, the overall top-performing systems from the remaining pool (both constrained and unconstrained) were added until a total of 18 systems was reached.
Limitations
A key limitation of our evaluation is that some models have been optimized for the very metrics we employ in AutoRank, either during training or at inference time Freitag et al. (2022a); Finkelstein and Freitag (2024). This can lead to artificially inflated scores that do not accurately reflect a model’s true capabilities Kovacs et al. (2024). To mitigate this issue, we aggregate the assessments from multiple learned metrics and LLM-as-a-judge approaches. However, even this strategy has shortcomings. First, scores from different learned metrics often exhibit high correlation among themselves. Second, LLM-as-a-judge approaches, including Gemba-ESA, may themselves have been leveraged to optimize machine translation models.
Another limitation is that we use automatic metrics to evaluate entire paragraphs, whereas their reliability is typically established at the sentence level. Additionally, learned metrics struggle with low-resource languages, such as English-to-Bhojpuri and English-to-Maasai. For these cases, we rely on chrF++ instead. However, chrF++ is a surface-level metric that, like BLEU, has repeatedly been shown to correlate poorly with human judgments (Kocmi et al., 2021; Freitag et al., 2022b, 2023).
Furthermore, our automatic evaluation is conducted at the paragraph level, without incorporating document-level context. This may lead to inflated scores for systems that translate the dataset paragraph by paragraph, disregarding dependencies and coherence across paragraphs.
The LLM-as-a-judge approach also depends on the language performance of the underlying LLMs. For our evaluation, we selected two top-performing multilingual systems: GPT-4.1 and Command A. Command A officially supports only 23 languages (Cohere Team, 2025), while the set of languages supported by GPT-4.1 is not publicly documented. Nevertheless, as both metrics correlate well across all languages and show strong agreement with other evaluation metrics, we retained them as judges for all 30 language pairs.
Finally, using automatically generated speech recognition transcripts as source text in the speech domain introduces additional noise, as the evaluation metrics are unlikely to be robust to ASR errors. Consequently, systems that handle the speech domain well may receive lower scores if their outputs diverge from the ASR transcript, even when their translations are correct.
Given these challenges, along with the well-documented biases and limitations of automatic metrics (Karpinska et al., 2022; Moghe et al., 2024), human evaluation remains indispensable. Accordingly, results from human assessments will supersede the automatic rankings presented here.
Acknowledgement
This report would not have been possible without the partnership with Árni Magnússon Institute for Icelandic Studies, Charles University, Cohere, Custom.MT, Dubformer, Gates Foundation, Google, Institute of the Estonian Language, Microsoft, NTT, Toloka, University of Tartu, University of Tokyo. Furthermore, we are grateful to Toshiaki Nakazawa.
English-Egyptian Arabic
System Name LP Supported Params. (B) Humeval? AutoRank $\downarrow$ CometKiwi-XL $\uparrow$ GEMBA-ESA-CMDA $\uparrow$ GEMBA-ESA-GPT4.1 $\uparrow$ MetricX-24-Hybrid-XL $\uparrow$ XCOMET-XL $\uparrow$
Shy-hunyuan-MT 7 1.0 0.658 76.3 75.0 -5.7 0.388
Wenyiil 14 2.5 0.65 79.2 73.3 -6.4 0.337
Algharb 14 2.6 0.645 80.0 73.9 -6.5 0.328
GemTrans 27 3.4 0.644 73.0 69.6 -6.0 0.345
CommandA-WMT 111 4.0 0.621 77.8 75.4 -7.0 0.311
UvA-MT 12 4.1 0.637 74.4 73.4 -7.1 0.325
Yolu 14 5.4 0.658 67.8 63.9 -6.6 0.323
$\blacktriangle$ Gemini-2.5-Pro ? 5.6 0.552 79.5 84.5 -7.6 0.267
$\blacktriangle$ ONLINE-B ? 6.4 0.627 70.4 67.4 -7.1 0.288
$\blacktriangle$ GPT-4.1 ? 6.5 0.534 78.4 84.1 -7.8 0.265
$\blacktriangle$ DeepSeek-V3 ? 671 6.9 0.573 74.2 75.7 -7.7 0.273
$\blacktriangle$ Mistral-Medium ? 7.5 0.586 71.7 71.0 -7.8 0.274
$\blacktriangle$ Claude-4 ? 7.6 0.552 76.5 80.0 -8.5 0.246
SRPOL 12 7.9 0.641 65.7 61.7 -7.8 0.286
$\blacktriangle$ CommandA 111 8.3 0.533 75.8 80.0 -8.5 0.238
$\blacktriangle$ AyaExpanse-32B 32 8.4 0.585 70.7 68.8 -8.1 0.261
$\blacktriangle$ ONLINE-W ? ? 9.0 0.607 67.7 64.0 -8.2 0.258
$\blacktriangle$ AyaExpanse-8B 8 9.7 0.596 66.1 61.6 -8.2 0.259
$\blacktriangle$ Qwen3-235B 235 10.7 0.571 66.1 64.1 -8.7 0.247
$\blacktriangle$ Gemma-3-27B 27 10.7 0.549 64.8 63.3 -8.6 0.281
$\blacktriangle$ EuroLLM-22B-pre.[M] 22 10.7 0.592 64.0 60.5 -8.5 0.246
IRB-MT 12 10.8 0.532 69.0 67.5 -8.5 0.236
$\blacktriangle$ Llama-4-Maverick 400 11.1 0.526 67.9 70.0 -8.8 0.234
IR-MultiagentMT ? 11.3 0.543 66.0 64.2 -8.7 0.247
$\blacktriangle$ CommandR7B 7 11.3 0.588 62.7 59.0 -8.8 0.248
$\blacktriangle$ Gemma-3-12B 12 11.7 0.529 67.9 67.6 -9.0 0.22
$\blacktriangle$ EuroLLM-9B[M] 9 14.0 0.548 58.7 54.5 -9.3 0.233
$\blacktriangle$ TowerPlus-72B[M] 72 15.5 0.534 58.2 54.0 -10.5 0.224
TranssionTranslate ? ? 15.8 0.501 59.0 57.4 -9.9 0.2
TranssionMT 1 16.9 0.488 58.5 56.1 -10.4 0.194
$\blacktriangle$ NLLB 1 18.0 0.499 53.9 51.2 -10.8 0.201
SalamandraTA 8 20.1 0.492 50.0 44.4 -11.4 0.195
$\blacktriangle$ ONLINE-G ? 22.6 0.445 53.5 48.3 -13.5 0.152
$\blacktriangle$ Llama-3.1-8B 8 22.8 0.458 45.5 41.8 -12.3 0.18
$\blacktriangle$ Qwen2.5-7B 7 24.0 0.436 44.5 39.3 -12.6 0.176
$\blacktriangle$ TowerPlus-9B[M] 9 31.9 0.337 31.1 26.9 -15.2 0.162
$\blacktriangle$ Mistral-7B 7 37.0 0.262 27.9 23.2 -18.4 0.157
English-Bhojpuri
System Name LP Supported Params. (B) Humeval? AutoRank $\downarrow$ chrF++ $\uparrow$
$\blacktriangle$ Gemini-2.5-Pro ? 1.0 40.6
Wenyiil 14 2.5 38.9
Algharb 14 2.8 38.6
$\blacktriangle$ ONLINE-B ? 4.1 37.1
TranssionTranslate ? ? 4.4 36.9
$\blacktriangle$ Claude-4 ? ? 4.5 36.7
$\blacktriangle$ DeepSeek-V3 ? 671 5.1 36.0
$\blacktriangle$ GPT-4.1 ? ? 5.5 35.6
Yolu 14 5.6 35.4
TranssionMT 1 6.2 34.8
$\blacktriangle$ Llama-4-Maverick 400 6.5 34.4
$\blacktriangle$ CommandA 111 6.5 34.4
$\blacktriangle$ NLLB 1 6.6 34.3
$\blacktriangle$ Gemma-3-27B ? 27 8.3 32.4
CommandA-WMT 111 8.8 31.8
COILD-BHO 7 8.9 31.8
$\blacktriangle$ Mistral-Medium ? ? 9.0 31.6
$\blacktriangle$ Qwen3-235B 235 11.1 29.2
IRB-MT 12 11.4 28.9
$\blacktriangle$ AyaExpanse-32B 32 11.4 28.9
Shy-hunyuan-MT 7 11.5 28.8
GemTrans 27 11.9 28.3
SalamandraTA 8 12.1 28.2
$\blacktriangle$ Gemma-3-12B ? 12 12.3 27.9
$\blacktriangle$ TowerPlus-9B[M] 9 12.7 27.4
$\blacktriangle$ TowerPlus-72B[M] 72 12.8 27.3
$\blacktriangle$ EuroLLM-22B-pre.[M] 22 13.6 26.4
$\blacktriangle$ EuroLLM-9B[M] 9 14.7 25.2
IR-MultiagentMT ? 15.9 23.9
$\blacktriangle$ CommandR7B 7 16.7 22.9
$\blacktriangle$ AyaExpanse-8B 8 16.7 22.9
$\blacktriangle$ Qwen2.5-7B ? 7 17.7 21.8
$\blacktriangle$ Mistral-7B 7 20.9 18.2
UvA-MT 12 28.4 9.7
$\blacktriangle$ Llama-3.1-8B 8 35.0 2.3
English-Czech
System Name LP Supported Params. (B) Humeval? AutoRank $\downarrow$ CometKiwi-XL $\uparrow$ GEMBA-ESA-CMDA $\uparrow$ GEMBA-ESA-GPT4.1 $\uparrow$ MetricX-24-Hybrid-XL $\uparrow$ XCOMET-XL $\uparrow$
Shy-hunyuan-MT 7 1.0 0.658 83.7 89.4 -5.5 0.639
$\blacktriangle$ Gemini-2.5-Pro ? 3.4 0.633 83.8 91.5 -6.2 0.574
CommandA-WMT 111 3.5 0.645 81.3 86.2 -6.0 0.594
$\blacktriangle$ GPT-4.1 ? 3.9 0.63 84.2 89.7 -6.6 0.576
Wenyiil 14 4.4 0.645 79.4 86.3 -6.4 0.586
$\blacktriangle$ DeepSeek-V3 ? 671 5.0 0.628 81.4 87.0 -6.5 0.565
GemTrans 27 5.0 0.636 76.6 81.8 -5.8 0.596
Algharb 14 6.2 0.627 79.4 85.0 -6.9 0.552
Yolu 14 6.3 0.651 74.6 78.6 -6.5 0.582
UvA-MT 12 6.4 0.637 77.3 82.9 -6.9 0.562
$\blacktriangle$ Mistral-Medium ? ? 7.0 0.621 78.4 84.4 -7.1 0.547
SRPOL 12 8.6 0.641 72.9 76.2 -7.3 0.552
$\blacktriangle$ CommandA 111 8.6 0.609 78.2 82.5 -7.6 0.524
Laniqo 9 8.6 0.643 67.3 69.9 -6.5 0.608
$\blacktriangle$ Claude-4 ? ? 8.8 0.606 78.6 83.0 -7.9 0.522
$\blacktriangle$ Gemma-3-27B 27 9.0 0.606 76.9 81.5 -7.5 0.523
$\blacktriangle$ ONLINE-B ? 10.2 0.612 73.1 77.0 -7.4 0.513
$\blacktriangle$ AyaExpanse-32B 32 10.2 0.604 74.2 78.9 -7.8 0.519
SalamandraTA 8 10.3 0.624 70.1 74.5 -7.3 0.528
$\blacktriangle$ Llama-4-Maverick 400 11.1 0.595 75.3 79.7 -8.3 0.494
$\blacktriangle$ ONLINE-W ? ? 11.2 0.602 74.5 77.9 -8.3 0.495
$\blacktriangle$ TowerPlus-9B[M] 9 11.4 0.605 72.0 75.8 -7.9 0.505
$\blacktriangle$ Qwen3-235B 235 11.5 0.599 71.8 76.0 -7.8 0.505
CUNI-MH-v2 9 11.9 0.609 69.2 73.4 -7.9 0.517
$\blacktriangle$ EuroLLM-22B-pre.[M] 22 12.5 0.593 72.2 75.0 -8.4 0.488
IRB-MT 12 12.6 0.591 71.0 73.6 -7.8 0.484
$\blacktriangle$ TowerPlus-72B[M] 72 12.9 0.592 70.8 74.9 -8.4 0.485
TranssionTranslate ? ? 13.2 0.597 68.5 72.0 -7.8 0.48
$\blacktriangle$ Gemma-3-12B 12 13.4 0.583 71.6 74.1 -8.5 0.48
CUNI-SFT 9 15.9 0.575 66.9 68.2 -8.9 0.468
$\blacktriangle$ AyaExpanse-8B 8 16.0 0.572 67.1 67.9 -8.7 0.457
CUNI-DocTransformer <1 17.5 0.558 68.7 71.1 -10.0 0.425
IR-MultiagentMT ? 17.7 0.546 66.5 68.7 -9.3 0.442
$\blacktriangle$ EuroLLM-9B[M] 9 18.9 0.527 63.1 63.7 -9.0 0.466
$\blacktriangle$ NLLB 1 25.5 0.485 55.7 57.3 -10.8 0.392
$\blacktriangle$ CommandR7B 7 28.0 0.457 58.5 51.3 -11.6 0.369
$\blacktriangle$ ONLINE-G ? 28.7 0.472 58.1 58.0 -12.8 0.313
$\blacktriangle$ Llama-3.1-8B 8 28.9 0.48 55.7 52.0 -12.1 0.317
$\blacktriangle$ Qwen2.5-7B ? 7 37.5 0.41 46.2 43.8 -14.2 0.239
$\blacktriangle$ Mistral-7B 7 40.8 0.374 45.8 41.2 -15.5 0.207
ctpc_nlp ? ? 41.1 0.369 43.3 39.8 -14.8 0.207
TranssionMT 1 42.0 0.364 45.4 45.4 -16.7 0.196
English-Estonian
System Name LP Supported Params. (B) Humeval? AutoRank $\downarrow$ CometKiwi-XL $\uparrow$ GEMBA-ESA-CMDA $\uparrow$ GEMBA-ESA-GPT4.1 $\uparrow$ MetricX-24-Hybrid-XL $\uparrow$ XCOMET-XL $\uparrow$
Shy-hunyuan-MT 7 1.0 0.72 78.8 87.8 -7.3 0.628
$\blacktriangle$ Gemini-2.5-Pro ? 2.5 0.7 74.1 90.7 -8.0 0.59
Wenyiil 14 2.6 0.708 74.4 86.0 -8.0 0.599
$\blacktriangle$ GPT-4.1 ? 3.0 0.695 75.2 87.9 -8.6 0.577
Yolu 14 3.7 0.72 72.1 77.4 -8.3 0.587
Algharb 14 3.8 0.692 73.6 84.1 -8.7 0.558
GemTrans 27 4.9 0.689 70.8 74.3 -8.3 0.558
Laniqo 9 5.1 0.711 67.2 68.1 -8.2 0.602
SRPOL 12 5.5 0.705 70.5 74.2 -9.7 0.538
UvA-MT 12 5.8 0.696 71.9 72.6 -10.0 0.531
$\blacktriangle$ ONLINE-B ? 5.8 0.678 69.9 76.5 -9.2 0.521
CommandA-WMT 111 5.9 0.689 71.6 71.8 -9.7 0.527
SalamandraTA 8 6.1 0.695 68.4 71.5 -9.3 0.532
$\blacktriangle$ Claude-4 ? ? 6.3 0.673 71.4 77.3 -10.6 0.505
TranssionTranslate ? ? 7.2 0.669 66.1 73.2 -9.5 0.501
$\blacktriangle$ Gemma-3-27B 27 7.4 0.662 70.2 71.8 -10.8 0.491
$\blacktriangle$ EuroLLM-22B-pre.[M] 22 7.9 0.654 68.6 72.2 -10.8 0.479
$\blacktriangle$ Llama-4-Maverick 400 8.0 0.655 69.0 71.9 -11.1 0.474
$\blacktriangle$ ONLINE-W ? ? 8.6 0.654 67.9 70.3 -11.6 0.471
$\blacktriangle$ DeepSeek-V3 ? 671 10.1 0.613 64.0 66.5 -11.4 0.468
IRB-MT 12 11.1 0.609 65.6 60.5 -11.8 0.413
IR-MultiagentMT ? 11.1 0.605 64.5 62.7 -11.9 0.423
$\blacktriangle$ Gemma-3-12B 12 12.1 0.597 65.6 59.4 -13.0 0.387
$\blacktriangle$ EuroLLM-9B[M] 9 13.5 0.522 57.3 55.0 -11.0 0.463
$\blacktriangle$ Mistral-Medium ? ? 13.9 0.574 59.9 54.8 -13.5 0.4
$\blacktriangle$ Qwen3-235B 235 14.1 0.576 62.6 54.1 -13.7 0.349
$\blacktriangle$ CommandA 111 15.9 0.546 63.9 48.4 -15.4 0.316
$\blacktriangle$ NLLB 1 16.1 0.528 56.6 53.4 -14.2 0.35
$\blacktriangle$ ONLINE-G ? 16.9 0.532 57.2 55.3 -15.6 0.297
$\blacktriangle$ TowerPlus-72B[M] 72 20.2 0.491 54.7 40.3 -17.0 0.254
TranssionMT 1 23.7 0.436 46.6 43.1 -19.1 0.176
$\blacktriangle$ Llama-3.1-8B 8 24.5 0.424 47.8 33.7 -18.7 0.166
$\blacktriangle$ TowerPlus-9B[M] 9 27.2 0.403 42.1 13.2 -19.0 0.19
$\blacktriangle$ AyaExpanse-32B 32 32.3 0.284 33.4 20.2 -23.2 0.135
$\blacktriangle$ Qwen2.5-7B ? 7 33.6 0.273 27.6 17.8 -23.6 0.144
$\blacktriangle$ CommandR7B 7 35.8 0.169 23.4 9.2 -22.6 0.193
$\blacktriangle$ Mistral-7B 7 37.4 0.182 18.1 11.4 -24.5 0.151
$\blacktriangle$ AyaExpanse-8B 8 38.0 0.151 17.4 10.1 -24.7 0.171
English-Icelandic
System Name LP Supported Params. (B) Humeval? AutoRank $\downarrow$ CometKiwi-XL $\uparrow$ GEMBA-ESA-CMDA $\uparrow$ GEMBA-ESA-GPT4.1 $\uparrow$ MetricX-24-Hybrid-XL $\uparrow$ XCOMET-XL $\uparrow$
Shy-hunyuan-MT 7 1.0 0.663 71.6 83.9 -7.5 0.543
$\blacktriangle$ Gemini-2.5-Pro ? 1.8 0.647 69.2 87.6 -7.7 0.512
$\blacktriangle$ GPT-4.1 ? 1.9 0.653 70.2 84.5 -8.3 0.516
Erlendur ? 2.2 0.646 69.5 85.1 -8.2 0.506
$\blacktriangle$ TowerPlus-9B[M] 9 3.9 0.64 67.1 76.3 -8.8 0.471
$\blacktriangle$ ONLINE-B ? 4.4 0.636 66.1 73.5 -8.8 0.464
$\blacktriangle$ Claude-4 ? ? 5.2 0.628 67.5 73.8 -10.6 0.43
$\blacktriangle$ TowerPlus-72B[M] 72 5.7 0.621 66.7 67.7 -10.1 0.435
TranssionTranslate ? ? 5.8 0.625 63.2 68.9 -9.1 0.43
UvA-MT 12 6.8 0.627 68.1 59.1 -11.6 0.402
CommandA-WMT 111 6.8 0.619 68.0 57.4 -11.1 0.404
GemTrans 27 7.0 0.609 65.0 59.1 -9.7 0.401
AMI 3 7.4 0.627 59.6 58.1 -9.7 0.426
SalamandraTA 8 8.6 0.605 61.6 53.9 -11.0 0.386
$\blacktriangle$ Llama-4-Maverick 400 8.8 0.587 64.7 58.8 -12.3 0.357
$\blacktriangle$ Mistral-Medium ? ? 9.7 0.583 65.3 51.5 -13.0 0.337
$\blacktriangle$ Gemma-3-27B 27 9.7 0.572 62.2 54.9 -12.4 0.364
$\blacktriangle$ DeepSeek-V3 ? 671 10.5 0.547 58.0 56.6 -12.1 0.378
IRB-MT 12 11.9 0.542 61.2 47.2 -13.6 0.306
IR-MultiagentMT ? 12.1 0.53 60.0 51.3 -13.7 0.31
$\blacktriangle$ Qwen3-235B 235 13.5 0.525 60.5 41.5 -15.0 0.275
$\blacktriangle$ Gemma-3-12B 12 13.8 0.517 60.3 42.1 -15.4 0.268
$\blacktriangle$ NLLB 1 15.2 0.477 53.0 48.2 -15.0 0.27
$\blacktriangle$ ONLINE-G ? 15.8 0.477 53.4 49.2 -16.1 0.243
$\blacktriangle$ CommandA 111 16.2 0.475 59.0 37.4 -17.0 0.221
$\blacktriangle$ Llama-3.1-8B 8 24.8 0.323 42.7 24.6 -21.3 0.133
$\blacktriangle$ EuroLLM-9B[M] 9 25.5 0.303 32.9 9.2 -17.4 0.237
$\blacktriangle$ AyaExpanse-32B 32 28.0 0.275 35.2 18.4 -23.3 0.145
$\blacktriangle$ CommandR7B 7 30.3 0.2 23.4 9.1 -20.9 0.216
$\blacktriangle$ EuroLLM-22B-pre.[M] 22 30.8 0.206 26.5 13.7 -23.7 0.171
$\blacktriangle$ Mistral-7B 7 31.8 0.177 25.2 14.3 -24.3 0.17
$\blacktriangle$ Qwen2.5-7B ? 7 31.8 0.186 24.1 13.1 -24.3 0.174
$\blacktriangle$ AyaExpanse-8B 8 33.0 0.153 21.7 11.3 -24.6 0.177
English-Italian
System Name LP Supported Params. (B) Humeval? AutoRank $\downarrow$ GEMBA-ESA-CMDA $\uparrow$ GEMBA-ESA-GPT4.1 $\uparrow$ MetricX-24-Hybrid-XL $\uparrow$ XCOMET-XL $\uparrow$
Shy-hunyuan-MT 7 1.0 84.6 88.7 -4.7 0.62
CommandA-WMT 111 2.6 83.4 88.0 -4.8 0.59
$\blacktriangle$ Gemini-2.5-Pro ? 4.4 85.5 90.5 -5.6 0.537
$\blacktriangle$ GPT-4.1 ? 4.5 85.0 89.8 -5.8 0.553
GemTrans 27 5.2 78.2 83.5 -4.9 0.581
UvA-MT 12 5.3 78.9 84.6 -5.4 0.595
$\blacktriangle$ DeepSeek-V3 ? 671 6.1 81.9 87.9 -5.9 0.543
$\blacktriangle$ Mistral-Medium ? ? 7.1 79.9 86.4 -6.0 0.544
$\blacktriangle$ Qwen3-235B 235 7.2 80.1 84.9 -5.8 0.541
Laniqo 9 7.6 70.5 75.3 -4.9 0.63
$\blacktriangle$ Claude-4 ? 8.4 81.7 85.2 -6.4 0.52
$\blacktriangle$ CommandA 111 8.5 79.4 83.7 -6.2 0.537
$\blacktriangle$ ONLINE-B ? 9.4 76.7 78.6 -5.6 0.53
$\blacktriangle$ TowerPlus-72B[M] 72 9.4 76.2 81.8 -6.1 0.539
$\blacktriangle$ AyaExpanse-32B 32 10.1 75.7 80.9 -6.1 0.527
$\blacktriangle$ ONLINE-W ? ? 10.1 74.6 81.1 -6.0 0.531
IRB-MT 12 10.2 73.8 79.8 -5.7 0.523
SalamandraTA 8 10.3 71.9 76.9 -5.8 0.561
$\blacktriangle$ EuroLLM-22B-pre.[M] 22 11.0 74.2 79.8 -6.4 0.53
TranssionTranslate ? ? 11.0 72.9 77.1 -5.7 0.523
$\blacktriangle$ TowerPlus-9B[M] 9 11.3 73.5 78.6 -6.2 0.526
$\blacktriangle$ Gemma-3-27B 27 12.6 73.4 78.3 -6.7 0.513
IR-MultiagentMT ? 13.6 73.0 77.0 -6.8 0.499
$\blacktriangle$ AyaExpanse-8B 8 14.9 69.5 73.9 -6.7 0.502
$\blacktriangle$ EuroLLM-9B[M] 9 15.2 68.3 73.5 -6.8 0.509
$\blacktriangle$ Gemma-3-12B 12 15.5 69.7 74.7 -7.1 0.494
$\blacktriangle$ Llama-4-Maverick 400 18.0 67.0 71.9 -7.5 0.479
$\blacktriangle$ CommandR7B 7 18.0 67.3 69.4 -7.4 0.486
$\blacktriangle$ Llama-3.1-8B 8 22.8 61.8 64.1 -8.1 0.449
$\blacktriangle$ Qwen2.5-7B 7 23.5 60.8 61.5 -7.8 0.44
$\blacktriangle$ NLLB 1 27.1 58.5 61.6 -9.3 0.421
$\blacktriangle$ ONLINE-G ? 30.0 58.7 60.6 -9.9 0.368
$\blacktriangle$ Mistral-7B 7 33.0 53.5 52.1 -9.8 0.363
English-Japanese
System Name LP Supported Params. (B) Humeval? AutoRank $\downarrow$ CometKiwi-XL $\uparrow$ GEMBA-ESA-CMDA $\uparrow$ GEMBA-ESA-GPT4.1 $\uparrow$ MetricX-24-Hybrid-XL $\uparrow$ XCOMET-XL $\uparrow$
Shy-hunyuan-MT 7 1.0 0.687 82.2 89.6 -5.5 0.592
In2x ? 72 2.3 0.711 78.4 86.3 -5.9 0.575
$\blacktriangle$ Gemini-2.5-Pro ? 2.4 0.672 83.2 91.2 -5.7 0.55
$\blacktriangle$ GPT-4.1 ? 2.9 0.674 81.8 89.7 -5.9 0.558
Wenyiil 14 2.9 0.682 79.6 88.6 -5.7 0.553
KIKIS 18 3.1 0.678 80.2 85.4 -5.5 0.551
Algharb 14 3.2 0.678 80.8 89.2 -5.8 0.541
CommandA-WMT 111 3.6 0.694 76.5 85.5 -5.8 0.55
$\blacktriangle$ DeepSeek-V3 ? 671 4.6 0.667 80.5 87.7 -6.2 0.531
$\blacktriangle$ Mistral-Medium ? ? 5.4 0.675 77.6 86.2 -6.4 0.532
GemTrans 27 5.5 0.667 72.3 80.2 -5.5 0.553
$\blacktriangle$ Claude-4 ? 5.7 0.677 78.3 86.3 -6.5 0.516
Yolu 14 5.9 0.697 69.9 77.9 -5.9 0.541
$\blacktriangle$ ONLINE-B ? 6.1 0.684 72.4 80.0 -6.0 0.527
UvA-MT 12 6.4 0.691 72.5 82.0 -6.3 0.517
bb88 ? ? 7.2 0.674 74.6 82.6 -6.5 0.498
$\blacktriangle$ CommandA 111 7.3 0.674 74.8 82.9 -6.6 0.504
$\blacktriangle$ Qwen3-235B 235 7.3 0.667 74.3 83.2 -6.4 0.499
Systran 18 7.3 0.703 68.7 77.1 -6.5 0.523
$\blacktriangle$ Gemma-3-27B 27 7.9 0.666 74.3 82.2 -6.6 0.497
NTTSU 14 8.0 0.676 67.7 74.3 -5.6 0.498
$\blacktriangle$ TowerPlus-72B[M] 72 8.6 0.671 71.4 80.6 -6.8 0.499
$\blacktriangle$ Llama-4-Maverick 400 9.1 0.661 71.5 81.1 -6.8 0.487
Laniqo 9 9.3 0.677 66.1 70.1 -6.3 0.529
$\blacktriangle$ AyaExpanse-32B 32 9.8 0.662 70.9 78.5 -6.8 0.472
IRB-MT 12 10.3 0.643 70.0 77.9 -6.5 0.474
$\blacktriangle$ TowerPlus-9B[M] 9 10.4 0.665 68.7 76.3 -6.9 0.477
SRPOL 12 10.8 0.683 66.7 73.9 -7.1 0.472
TranssionTranslate ? ? 12.0 0.668 64.6 71.8 -6.9 0.459
$\blacktriangle$ Gemma-3-12B 12 13.6 0.623 64.7 73.8 -7.0 0.461
$\blacktriangle$ AyaExpanse-8B 8 15.6 0.632 62.1 69.4 -7.4 0.422
$\blacktriangle$ ONLINE-W ? ? 16.4 0.611 61.7 67.5 -7.3 0.432
SH 56 16.4 0.641 59.9 65.6 -7.5 0.419
$\blacktriangle$ EuroLLM-22B-pre.[M] 22 16.8 0.623 62.0 69.4 -7.9 0.425
$\blacktriangle$ CommandR7B 7 19.9 0.62 59.3 65.1 -8.6 0.379
IR-MultiagentMT ? 22.9 0.576 54.6 62.1 -8.6 0.373
$\blacktriangle$ EuroLLM-9B[M] 9 23.6 0.561 53.8 60.5 -8.6 0.391
$\blacktriangle$ Qwen2.5-7B 7 24.5 0.594 54.6 58.6 -9.2 0.338
$\blacktriangle$ Llama-3.1-8B 8 25.5 0.596 51.4 54.8 -9.0 0.32
SalamandraTA 8 26.4 0.603 51.8 53.1 -9.4 0.299
$\blacktriangle$ NLLB 1 37.9 0.479 42.7 46.8 -11.5 0.245
$\blacktriangle$ ONLINE-G ? 40.7 0.495 45.0 45.8 -13.2 0.207
$\blacktriangle$ Mistral-7B 7 43.0 0.462 39.2 40.2 -12.6 0.193
English-Korean
System Name LP Supported Params. (B) Humeval? AutoRank $\downarrow$ CometKiwi-XL $\uparrow$ GEMBA-ESA-CMDA $\uparrow$ GEMBA-ESA-GPT4.1 $\uparrow$ MetricX-24-Hybrid-XL $\uparrow$ XCOMET-XL $\uparrow$
Shy-hunyuan-MT 7 1.0 0.697 83.8 85.6 -4.9 0.624
$\blacktriangle$ Gemini-2.5-Pro ? 2.5 0.683 85.3 88.1 -5.6 0.571
CommandA-WMT 111 2.8 0.711 79.6 82.3 -5.6 0.584
$\blacktriangle$ GPT-4.1 ? 2.8 0.686 83.6 86.3 -5.7 0.581
Wenyiil 14 2.9 0.691 82.1 85.0 -5.6 0.576
Algharb 14 3.0 0.687 83.2 85.9 -5.7 0.565
UvA-MT 12 4.2 0.706 78.2 81.1 -6.0 0.554
$\blacktriangle$ Claude-4 ? 4.3 0.694 82.0 84.6 -6.3 0.536
GemTrans 27 4.9 0.677 76.7 78.8 -5.4 0.568
$\blacktriangle$ DeepSeek-V3 ? 671 5.0 0.681 79.5 83.7 -6.1 0.539
$\blacktriangle$ CommandA 111 5.8 0.692 77.8 80.9 -6.6 0.524
$\blacktriangle$ Mistral-Medium ? ? 6.0 0.684 77.2 79.4 -6.3 0.53
$\blacktriangle$ Qwen3-235B 235 6.3 0.678 77.1 80.0 -6.2 0.509
Yolu 14 6.8 0.701 70.1 73.0 -5.9 0.533
$\blacktriangle$ ONLINE-B ? 7.8 0.679 72.8 73.8 -6.2 0.504
IRB-MT 12 8.4 0.657 74.9 76.3 -6.4 0.489
$\blacktriangle$ TowerPlus-72B[M] 72 8.5 0.684 72.2 75.2 -6.8 0.487
$\blacktriangle$ AyaExpanse-32B 32 8.6 0.673 72.1 75.7 -6.7 0.493
$\blacktriangle$ Llama-4-Maverick 400 8.8 0.665 73.6 75.0 -6.7 0.487
Laniqo 9 8.9 0.689 64.9 66.2 -6.1 0.54
$\blacktriangle$ Gemma-3-12B 12 9.0 0.667 73.6 77.0 -7.0 0.474
$\blacktriangle$ TowerPlus-9B[M] 9 9.8 0.678 70.2 73.5 -7.2 0.472
$\blacktriangle$ ONLINE-W ? ? 10.4 0.674 67.7 69.4 -6.8 0.467
TranssionTranslate ? ? 12.0 0.675 64.0 65.0 -6.9 0.439
$\blacktriangle$ AyaExpanse-8B 8 12.7 0.657 64.6 67.7 -7.3 0.434
$\blacktriangle$ Gemma-3-27B 27 12.9 0.626 67.1 67.9 -7.4 0.477
$\blacktriangle$ EuroLLM-22B-pre.[M] 22 13.0 0.654 66.4 68.9 -7.7 0.422
IR-MultiagentMT ? 16.3 0.614 61.2 64.2 -8.1 0.41
$\blacktriangle$ CommandR7B 7 18.2 0.619 59.8 61.3 -8.8 0.364
$\blacktriangle$ EuroLLM-9B[M] 9 19.1 0.594 56.8 57.7 -8.3 0.39
SalamandraTA 8 22.8 0.624 50.6 50.3 -9.7 0.29
$\blacktriangle$ Llama-3.1-8B 8 24.8 0.586 50.7 50.8 -10.2 0.278
$\blacktriangle$ Qwen2.5-7B 7 25.0 0.568 47.9 48.7 -9.4 0.291
$\blacktriangle$ NLLB 1 28.1 0.549 42.8 44.3 -10.4 0.286
$\blacktriangle$ ONLINE-G ? 32.4 0.532 44.8 44.3 -12.8 0.187
$\blacktriangle$ Mistral-7B 7 36.0 0.478 37.8 39.3 -12.7 0.174
English-Maasai
System Name LP Supported Params. (B) Humeval? AutoRank $\downarrow$ chrF++ $\uparrow$
Shy-hunyuan-MT 7 1.0 27.7
$\blacktriangle$ Claude-4 ? ? 2.6 26.1
$\blacktriangle$ Qwen3-235B 235 3.0 25.6
$\blacktriangle$ Llama-4-Maverick 400 3.2 25.4
$\blacktriangle$ CommandR7B 7 4.3 24.3
$\blacktriangle$ TowerPlus-9B[M] 9 5.3 23.2
TranssionMT 1 5.9 22.6
$\blacktriangle$ Gemini-2.5-Pro ? ? 6.1 22.5
$\blacktriangle$ DeepSeek-V3 ? 671 6.2 22.4
CommandA-WMT 111 6.4 22.2
$\blacktriangle$ AyaExpanse-32B 32 7.1 21.4
$\blacktriangle$ CommandA 111 7.9 20.6
$\blacktriangle$ Llama-3.1-8B 8 8.1 20.4
$\blacktriangle$ EuroLLM-9B[M] 9 8.2 20.3
$\blacktriangle$ EuroLLM-22B-pre.[M] 22 8.2 20.3
$\blacktriangle$ AyaExpanse-8B 8 8.2 20.2
$\blacktriangle$ Qwen2.5-7B ? 7 8.6 19.9
$\blacktriangle$ TowerPlus-72B[M] 72 8.8 19.7
$\blacktriangle$ Gemma-3-12B ? 12 8.8 19.6
IR-MultiagentMT ? 9.0 19.5
IRB-MT 12 9.7 18.7
$\blacktriangle$ Mistral-7B 7 11.3 17.1
$\blacktriangle$ Gemma-3-27B ? 27 13.3 15.1
UvA-MT 12 14.7 13.6
$\blacktriangle$ GPT-4.1 ? ? 14.9 13.4
GemTrans 27 16.7 11.6
$\blacktriangle$ NLLB 1 27.0 0.9
English-Russian
System Name LP Supported Params. (B) Humeval? AutoRank $\downarrow$ CometKiwi-XL $\uparrow$ GEMBA-ESA-CMDA $\uparrow$ GEMBA-ESA-GPT4.1 $\uparrow$ MetricX-24-Hybrid-XL $\uparrow$ XCOMET-XL $\uparrow$
Shy-hunyuan-MT 7 1.0 0.657 84.3 85.9 -4.9 0.652
CommandA-WMT 111 4.2 0.656 81.3 80.5 -5.8 0.607
$\blacktriangle$ Gemini-2.5-Pro ? 4.3 0.634 85.9 87.8 -6.1 0.575
Yandex ? 4.4 0.638 81.2 80.6 -5.3 0.617
UvA-MT 12 4.5 0.662 78.6 80.5 -6.1 0.611
Wenyiil 14 4.7 0.644 82.5 84.1 -6.1 0.588
GemTrans 27 5.1 0.639 77.8 79.5 -5.3 0.617
Algharb 14 5.1 0.637 84.4 85.5 -6.4 0.573
$\blacktriangle$ GPT-4.1 ? 5.3 0.631 84.6 85.8 -6.5 0.577
$\blacktriangle$ DeepSeek-V3 ? 671 5.6 0.632 84.2 84.7 -6.4 0.57
Yolu 14 6.9 0.658 73.1 73.6 -6.0 0.596
$\blacktriangle$ Claude-4 ? 8.5 0.619 82.0 81.6 -7.5 0.548
$\blacktriangle$ Qwen3-235B 235 8.7 0.625 78.2 79.9 -6.9 0.543
$\blacktriangle$ Gemma-3-27B 27 8.7 0.626 78.9 79.6 -7.3 0.551
Laniqo 9 8.7 0.649 67.9 67.0 -6.0 0.622
RuZh ? 9 9.5 0.633 74.5 74.7 -7.0 0.558
IRB-MT 12 9.9 0.616 75.8 76.5 -6.7 0.541
SRPOL 12 10.5 0.647 71.8 71.6 -7.7 0.549
$\blacktriangle$ TowerPlus-72B[M] 72 10.5 0.624 75.7 75.2 -7.6 0.543
$\blacktriangle$ CommandA 111 11.0 0.618 76.9 76.3 -8.1 0.536
$\blacktriangle$ ONLINE-W ? ? 11.5 0.624 73.7 73.6 -7.9 0.534
DLUT_GTCOM 27 11.6 0.626 71.1 71.2 -7.3 0.537
$\blacktriangle$ TowerPlus-9B[M] 9 11.8 0.617 73.2 72.7 -7.5 0.533
$\blacktriangle$ Llama-4-Maverick 400 12.3 0.616 75.3 75.5 -8.5 0.513
SalamandraTA 8 12.3 0.632 69.6 68.2 -7.6 0.534
$\blacktriangle$ ONLINE-B ? 12.7 0.616 73.8 72.8 -8.2 0.517
TranssionTranslate ? ? 12.7 0.618 70.7 71.0 -7.5 0.52
$\blacktriangle$ AyaExpanse-32B 32 13.5 0.603 71.8 72.1 -7.9 0.517
$\blacktriangle$ ONLINE-G ? 14.2 0.613 67.8 66.6 -7.6 0.522
$\blacktriangle$ EuroLLM-22B-pre.[M] 22 14.4 0.606 70.9 71.3 -8.4 0.502
$\blacktriangle$ Gemma-3-12B 12 14.7 0.589 72.6 73.2 -8.4 0.503
$\blacktriangle$ AyaExpanse-8B 8 17.4 0.589 67.1 66.2 -8.7 0.48
IR-MultiagentMT ? 19.4 0.564 65.7 65.7 -8.9 0.467
$\blacktriangle$ EuroLLM-9B[M] 9 21.5 0.547 63.6 63.1 -9.5 0.471
$\blacktriangle$ Qwen2.5-7B 7 24.8 0.546 60.2 57.5 -10.2 0.411
$\blacktriangle$ Llama-3.1-8B 8 29.7 0.521 58.6 55.1 -12.6 0.372
$\blacktriangle$ NLLB 1 31.5 0.483 54.3 53.3 -11.7 0.389
TranssionMT 1 34.2 0.483 54.3 54.9 -13.7 0.332
$\blacktriangle$ Mistral-7B 7 39.0 0.45 52.4 46.3 -14.4 0.288
$\blacktriangle$ CommandR7B 7 40.0 0.41 52.1 39.9 -13.6 0.347
English-Serbian (Cyrilics)
System Name LP Supported Params. (B) Humeval? AutoRank $\downarrow$ CometKiwi-XL $\uparrow$ GEMBA-ESA-CMDA $\uparrow$ GEMBA-ESA-GPT4.1 $\uparrow$ MetricX-24-Hybrid-XL $\uparrow$ XCOMET-XL $\uparrow$
Shy-hunyuan-MT 7 1.0 0.687 76.6 83.3 -4.2 0.64
$\blacktriangle$ Gemini-2.5-Pro ? 3.0 0.663 74.6 87.2 -5.1 0.566
$\blacktriangle$ GPT-4.1 ? 3.4 0.655 74.4 83.4 -5.2 0.582
GemTrans 27 4.6 0.663 71.6 74.5 -4.9 0.554
UvA-MT 12 5.8 0.658 71.4 70.2 -4.5 0.46
$\blacktriangle$ ONLINE-B ? 6.1 0.644 71.0 75.2 -5.7 0.517
$\blacktriangle$ Claude-4 ? ? 6.8 0.628 72.6 77.4 -6.6 0.503
CommandA-WMT 111 7.0 0.641 71.8 67.3 -6.0 0.512
TranssionTranslate ? ? 8.0 0.631 67.3 70.9 -6.0 0.484
$\blacktriangle$ DeepSeek-V3 ? 671 8.6 0.603 68.0 72.0 -6.6 0.501
SalamandraTA 8 8.8 0.635 66.2 65.1 -6.2 0.48
DLUT_GTCOM 27 9.3 0.618 66.9 68.1 -6.6 0.463
IRB-MT 12 9.9 0.604 67.9 64.2 -6.5 0.435
$\blacktriangle$ Llama-4-Maverick 400 10.0 0.603 68.2 68.9 -7.3 0.444
$\blacktriangle$ Qwen3-235B 235 11.9 0.591 65.8 60.1 -7.6 0.425
$\blacktriangle$ Gemma-3-12B 12 12.1 0.583 65.7 61.8 -7.4 0.394
$\blacktriangle$ Gemma-3-27B 27 12.2 0.583 61.7 62.1 -7.4 0.444
CUNI-SFT 9 13.5 0.569 61.1 52.4 -5.8 0.328
IR-MultiagentMT ? 14.1 0.548 63.3 59.2 -8.1 0.386
$\blacktriangle$ ONLINE-G ? 14.5 0.566 58.8 56.2 -7.7 0.383
$\blacktriangle$ CommandA 111 17.6 0.527 62.9 50.8 -10.0 0.323
$\blacktriangle$ Llama-3.1-8B 8 19.4 0.489 53.9 44.2 -7.5 0.233
$\blacktriangle$ NLLB 1 19.8 0.468 53.5 50.3 -9.4 0.33
$\blacktriangle$ EuroLLM-22B-pre.[M] 22 20.6 0.469 53.6 41.4 -8.6 0.269
$\blacktriangle$ EuroLLM-9B[M] 9 22.4 0.454 51.5 37.4 -9.4 0.265
$\blacktriangle$ TowerPlus-72B[M] 72 26.0 0.424 51.6 36.9 -12.4 0.203
$\blacktriangle$ TowerPlus-9B[M] 9 26.7 0.368 43.7 29.2 -9.1 0.182
$\blacktriangle$ Mistral-7B 7 27.0 0.414 49.2 38.3 -13.0 0.207
$\blacktriangle$ AyaExpanse-8B 8 29.7 0.306 40.9 27.3 -10.1 0.157
$\blacktriangle$ CommandR7B 7 31.3 0.307 38.0 26.0 -11.6 0.171
$\blacktriangle$ AyaExpanse-32B 32 31.8 0.354 46.4 29.6 -15.2 0.142
$\blacktriangle$ Qwen2.5-7B ? 7 32.0 0.306 37.0 27.2 -11.9 0.144
English-Ukrainian
System Name LP Supported Params. (B) Humeval? AutoRank $\downarrow$ CometKiwi-XL $\uparrow$ GEMBA-ESA-CMDA $\uparrow$ GEMBA-ESA-GPT4.1 $\uparrow$ MetricX-24-Hybrid-XL $\uparrow$ XCOMET-XL $\uparrow$
Shy-hunyuan-MT 7 1.0 0.65 84.1 85.3 -5.0 0.662
$\blacktriangle$ Gemini-2.5-Pro ? 3.3 0.625 84.6 89.8 -6.3 0.59
Wenyiil 14 3.4 0.635 83.7 85.4 -6.2 0.597
$\blacktriangle$ GPT-4.1 ? 3.4 0.626 82.8 87.0 -6.2 0.611
CommandA-WMT 111 3.8 0.641 80.4 82.4 -6.0 0.599
Algharb 14 4.1 0.625 83.2 86.0 -6.5 0.586
UvA-MT 12 4.3 0.641 78.7 81.5 -6.3 0.6
GemTrans 27 4.5 0.628 78.0 80.1 -5.7 0.606
$\blacktriangle$ DeepSeek-V3 ? 671 4.9 0.619 81.7 84.0 -6.5 0.574
Yolu 14 5.9 0.643 73.4 74.4 -6.2 0.589
$\blacktriangle$ Mistral-Medium ? ? 5.9 0.617 79.8 82.1 -6.9 0.566
$\blacktriangle$ Claude-4 ? ? 6.9 0.604 81.1 82.6 -7.6 0.544
$\blacktriangle$ CommandA 111 7.3 0.61 78.1 79.8 -7.4 0.546
Laniqo 9 7.5 0.638 67.3 66.3 -6.3 0.613
IRB-MT 12 8.0 0.604 74.8 76.9 -6.9 0.539
SRPOL 12 8.2 0.631 70.9 72.9 -7.3 0.548
$\blacktriangle$ TowerPlus-9B[M] 9 8.6 0.603 73.4 75.2 -7.2 0.541
$\blacktriangle$ Llama-4-Maverick 400 8.6 0.603 76.3 78.1 -7.8 0.519
CGFOKUS 235 8.7 0.597 75.7 78.1 -7.4 0.513
$\blacktriangle$ ONLINE-B ? 8.8 0.609 73.2 73.3 -7.3 0.531
$\blacktriangle$ AyaExpanse-32B 32 9.1 0.6 73.9 75.0 -7.5 0.528
$\blacktriangle$ ONLINE-W ? ? 9.1 0.605 72.8 75.0 -7.5 0.527
$\blacktriangle$ Qwen3-235B 235 9.3 0.596 73.8 75.7 -7.5 0.515
SalamandraTA 8 9.9 0.613 68.3 68.6 -7.2 0.528
$\blacktriangle$ TowerPlus-72B[M] 72 10.2 0.592 72.4 73.8 -7.9 0.514
TranssionTranslate ? ? 10.7 0.594 69.1 71.3 -7.5 0.505
DLUT_GTCOM 27 11.0 0.592 69.8 71.4 -7.9 0.498
$\blacktriangle$ Gemma-3-27B 27 11.9 0.575 68.1 71.0 -8.1 0.51
$\blacktriangle$ EuroLLM-22B-pre.[M] 22 12.5 0.577 68.9 69.4 -8.6 0.492
$\blacktriangle$ AyaExpanse-8B 8 13.1 0.576 66.5 67.8 -8.4 0.477
CUNI-SFT 9 13.3 0.579 66.1 65.5 -8.5 0.484
$\blacktriangle$ ONLINE-G ? 13.7 0.575 64.2 65.0 -8.4 0.479
IR-MultiagentMT ? 14.0 0.555 67.3 67.3 -8.5 0.467
$\blacktriangle$ Gemma-3-12B 12 14.4 0.559 64.9 65.8 -8.6 0.473
$\blacktriangle$ EuroLLM-9B[M] 9 17.0 0.518 63.1 61.8 -9.0 0.459
$\blacktriangle$ NLLB 1 24.0 0.467 53.2 53.6 -11.2 0.368
$\blacktriangle$ Llama-3.1-8B 8 24.3 0.488 55.5 51.0 -11.9 0.331
TranssionMT 1 28.1 0.441 51.8 52.2 -13.5 0.286
$\blacktriangle$ CommandR7B 7 29.0 0.411 54.6 43.6 -13.2 0.323
$\blacktriangle$ Mistral-7B 7 29.3 0.428 52.2 46.2 -13.4 0.277
$\blacktriangle$ Qwen2.5-7B ? 7 36.6 0.362 41.8 36.0 -15.2 0.2
KYUoM ? <1 42.0 0.265 35.9 34.7 -16.6 0.201
English-Simplified Chinese
System Name LP Supported Params. (B) Humeval? AutoRank $\downarrow$ CometKiwi-XL $\uparrow$ GEMBA-ESA-CMDA $\uparrow$ GEMBA-ESA-GPT4.1 $\uparrow$ MetricX-24-Hybrid-XL $\uparrow$ XCOMET-XL $\uparrow$
Shy-hunyuan-MT 7 1.0 0.67 87.2 88.3 -4.0 0.576
Wenyiil 14 3.9 0.663 84.2 87.7 -5.0 0.52
$\blacktriangle$ Gemini-2.5-Pro ? 4.0 0.657 85.2 88.7 -4.9 0.512
Algharb 14 4.1 0.66 84.7 87.8 -5.0 0.515
$\blacktriangle$ GPT-4.1 ? 4.6 0.652 84.9 86.8 -5.0 0.512
$\blacktriangle$ Qwen3-235B 235 4.8 0.661 82.7 85.0 -5.0 0.513
Yolu 14 4.8 0.687 74.9 77.1 -4.6 0.542
GemTrans 27 4.9 0.658 77.0 80.2 -4.3 0.546
$\blacktriangle$ Mistral-Medium ? ? 4.9 0.658 82.4 84.9 -5.0 0.514
CommandA-WMT 111 5.6 0.665 78.9 81.5 -5.0 0.508
UvA-MT 12 6.3 0.671 76.8 81.0 -5.4 0.499
$\blacktriangle$ Claude-4 ? 7.0 0.649 80.4 82.8 -5.6 0.487
$\blacktriangle$ DeepSeek-V3 671 7.1 0.618 84.9 85.1 -5.2 0.473
$\blacktriangle$ Llama-4-Maverick 400 8.0 0.65 74.9 79.4 -5.5 0.489
$\blacktriangle$ ONLINE-B ? 8.2 0.656 73.0 74.7 -5.2 0.492
$\blacktriangle$ Gemma-3-27B 27 9.0 0.638 75.5 78.7 -5.8 0.475
Laniqo 9 9.1 0.665 65.6 67.4 -4.9 0.513
IRB-MT 12 9.3 0.633 73.7 77.5 -5.3 0.467
$\blacktriangle$ CommandA 111 9.4 0.645 76.5 76.8 -6.1 0.464
$\blacktriangle$ TowerPlus-72B[M] 72 9.8 0.645 73.3 76.6 -6.1 0.466
SRPOL 12 10.3 0.666 68.2 71.1 -6.0 0.461
RuZh ? 9 10.4 0.648 71.2 74.2 -5.9 0.454
$\blacktriangle$ Gemma-3-12B 12 10.6 0.636 73.4 76.6 -6.1 0.446
$\blacktriangle$ Qwen2.5-7B 7 11.5 0.625 70.6 73.6 -5.9 0.451
$\blacktriangle$ AyaExpanse-32B 32 11.6 0.631 70.9 74.6 -6.3 0.444
$\blacktriangle$ TowerPlus-9B[M] 9 11.9 0.634 69.9 71.7 -6.2 0.446
TranssionTranslate ? ? 12.7 0.638 66.9 70.1 -6.4 0.438
$\blacktriangle$ EuroLLM-22B-pre.[M] 22 12.8 0.627 68.9 71.5 -6.4 0.43
$\blacktriangle$ ONLINE-W ? ? 13.4 0.627 66.4 69.2 -6.5 0.437
$\blacktriangle$ AyaExpanse-8B 8 15.0 0.615 65.1 68.6 -6.7 0.403
$\blacktriangle$ EuroLLM-9B[M] 9 16.4 0.604 63.6 66.6 -6.9 0.394
IR-MultiagentMT ? 17.2 0.575 64.0 66.0 -6.6 0.399
SalamandraTA 8 17.9 0.618 59.5 59.2 -7.1 0.376
$\blacktriangle$ Llama-3.1-8B 8 18.4 0.594 61.6 62.8 -7.4 0.379
$\blacktriangle$ CommandR7B 7 18.5 0.595 63.1 65.0 -7.9 0.376
$\blacktriangle$ ONLINE-G ? 31.2 0.508 52.2 51.7 -11.1 0.256
$\blacktriangle$ Mistral-7B 7 32.0 0.5 47.6 46.7 -10.4 0.257
$\blacktriangle$ NLLB 1 38.0 0.441 44.4 45.6 -12.8 0.238
Czech-Ukrainian
System Name LP Supported Params. (B) Humeval? AutoRank $\downarrow$ CometKiwi-XL $\uparrow$ GEMBA-ESA-CMDA $\uparrow$ GEMBA-ESA-GPT4.1 $\uparrow$ MetricX-24-Hybrid-XL $\uparrow$ XCOMET-XL $\uparrow$
Shy-hunyuan-MT 7 1.0 0.601 79.1 85.3 -5.0 0.681
$\blacktriangle$ Gemini-2.5-Pro ? 1.0 0.582 81.4 89.5 -5.1 0.671
CommandA-WMT 111 1.3 0.593 80.3 84.3 -4.8 0.664
$\blacktriangle$ GPT-4.1 ? 1.3 0.592 80.4 89.0 -5.3 0.666
$\blacktriangle$ DeepSeek-V3 ? 671 3.2 0.578 79.3 84.3 -5.5 0.654
$\blacktriangle$ Claude-4 ? ? 3.6 0.587 78.8 85.6 -6.0 0.645
$\blacktriangle$ Mistral-Medium ? ? 4.1 0.58 77.9 83.5 -5.8 0.642
GemTrans 27 4.3 0.58 75.6 79.2 -5.2 0.645
$\blacktriangle$ CommandA 111 4.5 0.582 78.9 81.7 -6.0 0.637
$\blacktriangle$ Gemma-3-27B 27 4.9 0.581 77.3 81.7 -6.0 0.63
UvA-MT 12 5.0 0.597 74.7 79.1 -6.0 0.64
Wenyiil 14 5.3 0.585 75.6 79.1 -5.9 0.635
Yolu 14 5.9 0.606 72.1 73.8 -5.9 0.634
Algharb 14 7.1 0.572 74.0 79.5 -6.4 0.619
$\blacktriangle$ Llama-4-Maverick 400 7.3 0.574 75.5 80.3 -6.7 0.601
$\blacktriangle$ AyaExpanse-32B 32 7.4 0.57 73.4 76.1 -6.1 0.618
Laniqo 9 7.5 0.596 68.1 68.6 -5.9 0.645
SRPOL 12 7.6 0.6 71.4 73.3 -6.6 0.618
$\blacktriangle$ TowerPlus-9B[M] 9 7.7 0.57 74.0 76.7 -6.4 0.608
$\blacktriangle$ TowerPlus-72B[M] 72 8.7 0.567 72.7 75.7 -6.7 0.602
$\blacktriangle$ EuroLLM-22B-pre.[M] 22 8.7 0.566 72.8 74.7 -6.7 0.606
IRB-MT 12 8.9 0.559 72.4 74.8 -6.4 0.598
$\blacktriangle$ Gemma-3-12B 12 9.7 0.559 73.0 75.9 -6.9 0.583
$\blacktriangle$ Qwen3-235B 235 10.4 0.557 71.5 73.6 -6.9 0.582
IR-MultiagentMT ? 11.2 0.544 70.1 71.8 -6.7 0.579
$\blacktriangle$ ONLINE-B ? 11.5 0.542 69.1 69.8 -6.5 0.578
SalamandraTA 8 11.7 0.562 66.9 66.5 -6.7 0.583
CUNI-EdUKate-v1 9 12.5 0.555 67.2 67.4 -7.1 0.573
$\blacktriangle$ AyaExpanse-8B 8 13.3 0.54 66.9 66.7 -6.9 0.565
$\blacktriangle$ ONLINE-W ? ? 13.4 0.534 66.6 68.0 -6.9 0.564
TranssionTranslate ? ? 14.6 0.521 66.5 67.2 -7.0 0.541
$\blacktriangle$ EuroLLM-9B[M] 9 14.8 0.533 66.6 67.3 -7.6 0.545
DLUT_GTCOM 27 15.0 0.523 65.9 67.3 -7.3 0.54
CUNI-SFT 9 15.2 0.528 63.7 64.8 -7.2 0.552
$\blacktriangle$ ONLINE-G ? 24.0 0.471 58.0 55.3 -8.8 0.458
$\blacktriangle$ Llama-3.1-8B 8 25.3 0.493 58.2 53.8 -10.0 0.432
CUNI-Transformer <1 25.7 0.47 58.0 56.3 -10.1 0.449
$\blacktriangle$ CommandR7B 7 25.8 0.481 57.1 52.2 -10.2 0.467
$\blacktriangle$ NLLB 1 28.5 0.46 51.7 50.8 -10.3 0.439
TranssionMT 1 33.0 0.444 52.1 49.7 -12.1 0.371
$\blacktriangle$ Mistral-7B 7 33.5 0.443 52.7 47.2 -12.0 0.359
$\blacktriangle$ Qwen2.5-7B ? 7 42.0 0.382 45.6 40.2 -13.8 0.287
Czech-German
System Name LP Supported Params. (B) Humeval? AutoRank $\downarrow$ CometKiwi-XL $\uparrow$ GEMBA-ESA-CMDA $\uparrow$ GEMBA-ESA-GPT4.1 $\uparrow$ MetricX-24-Hybrid-XL $\uparrow$ XCOMET-XL $\uparrow$
Shy-hunyuan-MT 7 1.0 0.596 78.4 88.3 -3.6 0.653
CommandA-WMT 111 2.1 0.582 77.9 87.5 -3.2 0.634
$\blacktriangle$ GPT-4.1 ? 2.3 0.58 79.4 91.0 -3.7 0.634
$\blacktriangle$ Gemini-2.5-Pro ? 2.5 0.577 79.1 90.8 -3.6 0.633
$\blacktriangle$ DeepSeek-V3 ? 671 3.5 0.577 79.5 88.6 -3.8 0.624
$\blacktriangle$ Mistral-Medium ? 4.1 0.577 77.6 86.9 -3.8 0.627
$\blacktriangle$ CommandA 111 4.7 0.579 77.8 85.5 -4.0 0.624
$\blacktriangle$ Claude-4 ? 4.7 0.577 77.7 87.1 -3.9 0.618
GemTrans 27 6.2 0.569 75.0 82.1 -3.7 0.619
UvA-MT 12 6.8 0.584 74.2 82.3 -4.3 0.617
$\blacktriangle$ Gemma-3-27B 27 7.1 0.572 74.9 82.5 -4.1 0.612
$\blacktriangle$ Llama-4-Maverick 400 7.5 0.569 74.7 84.7 -4.2 0.604
$\blacktriangle$ AyaExpanse-32B 32 8.0 0.568 74.3 80.5 -4.1 0.606
Yolu 14 9.0 0.589 70.0 75.2 -4.4 0.613
$\blacktriangle$ TowerPlus-72B[M] 72 9.3 0.572 73.1 78.5 -4.4 0.6
$\blacktriangle$ Qwen3-235B 235 9.3 0.565 73.1 80.9 -4.2 0.594
Laniqo 9 10.1 0.587 67.5 70.3 -4.2 0.619
$\blacktriangle$ TowerPlus-9B[M] 9 10.1 0.568 71.7 77.4 -4.4 0.599
Wenyiil 14 10.7 0.559 71.1 77.9 -4.3 0.597
SRPOL 12 10.8 0.593 69.2 73.2 -4.7 0.591
$\blacktriangle$ EuroLLM-22B-pre.[M] 22 11.0 0.567 70.7 77.4 -4.6 0.596
$\blacktriangle$ Gemma-3-12B 12 11.2 0.561 71.9 77.5 -4.6 0.592
$\blacktriangle$ ONLINE-B ? 11.7 0.555 69.3 74.4 -4.1 0.597
IRB-MT 12 12.1 0.557 70.6 75.4 -4.5 0.588
Algharb 14 12.9 0.551 70.8 77.1 -4.7 0.58
IR-MultiagentMT ? 13.0 0.559 68.0 75.3 -4.7 0.592
CUNI-MH-v2 9 13.8 0.562 68.2 72.5 -4.7 0.577
SalamandraTA 8 15.3 0.554 65.8 69.5 -4.6 0.574
$\blacktriangle$ AyaExpanse-8B 8 15.4 0.555 66.4 70.9 -4.7 0.564
TranssionTranslate ? ? 16.6 0.538 67.0 71.1 -4.7 0.56
$\blacktriangle$ ONLINE-W ? ? 16.7 0.542 67.0 71.1 -4.9 0.56
DLUT_GTCOM 27 17.4 0.537 66.6 70.5 -4.8 0.553
$\blacktriangle$ CommandR7B 7 17.9 0.545 65.8 68.9 -5.1 0.556
$\blacktriangle$ EuroLLM-9B[M] 9 22.4 0.531 57.1 61.1 -5.6 0.579
$\blacktriangle$ Llama-3.1-8B 8 25.3 0.524 59.6 61.6 -5.8 0.508
$\blacktriangle$ ONLINE-G ? 32.1 0.492 58.1 58.4 -6.9 0.47
$\blacktriangle$ Qwen2.5-7B 7 32.1 0.503 54.0 54.5 -6.6 0.476
$\blacktriangle$ NLLB 1 33.4 0.5 52.5 54.2 -6.9 0.479
$\blacktriangle$ Mistral-7B 7 36.4 0.492 53.3 51.1 -7.1 0.434
TranssionMT 1 40.0 0.473 51.2 52.4 -7.9 0.425
Japanese-Simplified Chinese
System Name LP Supported Params. (B) Humeval? AutoRank $\downarrow$ CometKiwi-XL $\uparrow$ GEMBA-ESA-CMDA $\uparrow$ GEMBA-ESA-GPT4.1 $\uparrow$ MetricX-24-Hybrid-XL $\uparrow$ XCOMET-XL $\uparrow$
Shy-hunyuan-MT 7 1.0 0.577 85.1 85.5 -4.2 0.629
In2x ? 72 3.0 0.624 77.0 77.7 -4.7 0.618
$\blacktriangle$ Gemini-2.5-Pro ? 3.2 0.549 84.8 84.8 -4.6 0.596
Kaze-MT 72 3.8 0.569 81.5 81.8 -4.8 0.605
Algharb 14 4.2 0.547 83.5 84.1 -4.8 0.583
$\blacktriangle$ GPT-4.1 ? 4.4 0.549 83.8 84.7 -5.1 0.582
Wenyiil 14 4.5 0.555 81.4 81.9 -4.8 0.591
CommandA-WMT 111 5.1 0.558 80.2 79.7 -4.7 0.575
NTTSU 14 5.8 0.563 77.5 74.8 -4.6 0.577
bb88 ? ? 6.1 0.551 80.1 78.9 -5.2 0.573
$\blacktriangle$ Claude-4 ? 6.2 0.545 82.9 83.7 -5.6 0.556
$\blacktriangle$ DeepSeek-V3 671 6.3 0.534 82.9 80.9 -5.1 0.552
$\blacktriangle$ Mistral-Medium ? ? 6.4 0.546 81.1 81.1 -5.4 0.558
GemTrans 27 6.5 0.556 76.0 74.9 -4.8 0.579
Yolu 14 6.9 0.578 74.6 73.6 -5.0 0.565
$\blacktriangle$ Qwen3-235B 235 7.5 0.549 78.4 77.0 -5.4 0.555
$\blacktriangle$ CommandA 111 7.6 0.54 79.4 77.6 -5.5 0.556
UvA-MT 12 8.3 0.564 73.9 75.2 -5.6 0.561
$\blacktriangle$ TowerPlus-72B[M] 72 9.7 0.537 76.5 75.0 -5.9 0.536
$\blacktriangle$ AyaExpanse-32B 32 10.7 0.537 73.2 72.0 -5.8 0.521
Laniqo 9 11.1 0.579 63.1 62.1 -5.4 0.557
$\blacktriangle$ TowerPlus-9B[M] 9 11.2 0.535 71.9 69.8 -5.8 0.523
IRB-MT 12 12.1 0.521 72.2 70.4 -6.0 0.509
$\blacktriangle$ Gemma-3-27B 27 12.8 0.526 70.4 70.2 -6.2 0.503
$\blacktriangle$ Llama-4-Maverick 400 13.1 0.524 71.5 66.1 -6.3 0.518
$\blacktriangle$ Qwen2.5-7B 7 13.6 0.524 68.9 67.4 -6.3 0.502
IR-MultiagentMT ? 13.7 0.523 67.8 68.5 -6.2 0.492
SRPOL 12 13.8 0.56 63.8 62.5 -6.4 0.522
$\blacktriangle$ EuroLLM-22B-pre.[M] 22 14.7 0.521 66.4 66.2 -6.3 0.486
$\blacktriangle$ AyaExpanse-8B 8 15.5 0.518 65.6 64.4 -6.4 0.472
$\blacktriangle$ ONLINE-B ? 16.2 0.499 63.7 63.2 -6.2 0.472
$\blacktriangle$ Gemma-3-12B 12 17.1 0.509 65.0 64.1 -7.1 0.465
$\blacktriangle$ CommandR7B 7 18.4 0.496 59.8 58.5 -6.9 0.486
TranssionTranslate ? ? 18.8 0.488 59.9 60.6 -6.7 0.45
$\blacktriangle$ Llama-3.1-8B 8 20.2 0.507 58.8 57.3 -7.2 0.423
$\blacktriangle$ EuroLLM-9B[M] 9 20.8 0.479 59.4 57.2 -7.6 0.461
$\blacktriangle$ ONLINE-W ? ? 25.2 0.456 52.3 52.9 -7.9 0.387
$\blacktriangle$ Mistral-7B 7 32.8 0.445 42.9 43.4 -9.8 0.317
SalamandraTA 8 33.1 0.426 36.5 38.0 -8.6 0.328
$\blacktriangle$ ONLINE-G ? 40.8 0.352 39.5 39.8 -12.1 0.28
$\blacktriangle$ NLLB 1 41.0 0.371 35.5 35.8 -12.1 0.303
English-Bengali
System Name LP Supported Params. (B) AutoRank $\downarrow$ GEMBA-ESA-CMDA $\uparrow$ GEMBA-ESA-GPT4.1 $\uparrow$ MetricX-24-Hybrid-XL $\uparrow$ XCOMET-XL $\uparrow$
Shy-hunyuan-MT 7 1.0 67.9 83.2 -4.8 0.449
$\blacktriangle$ Gemini-2.5-Pro ? 2.2 66.5 86.6 -5.2 0.382
$\blacktriangle$ GPT-4.1 ? 3.2 66.9 81.6 -5.9 0.373
GemTrans 27 3.5 64.3 75.1 -5.0 0.374
$\blacktriangle$ Mistral-Medium ? ? 3.9 65.5 78.0 -6.0 0.366
$\blacktriangle$ Claude-4 ? 4.0 65.6 80.6 -6.1 0.348
UvA-MT ? 12 4.2 64.1 75.0 -6.1 0.381
$\blacktriangle$ DeepSeek-V3 ? 671 4.3 63.7 77.7 -6.1 0.364
IRB-MT 12 5.1 62.9 72.7 -6.0 0.34
CommandA-WMT 111 5.3 63.5 69.4 -6.2 0.345
$\blacktriangle$ Llama-4-Maverick 400 5.5 63.9 73.9 -6.3 0.315
$\blacktriangle$ Qwen3-235B 235 6.0 62.7 71.2 -6.4 0.313
$\blacktriangle$ ONLINE-B ? 7.1 59.5 65.9 -6.4 0.304
TranssionTranslate ? ? 7.3 59.4 63.9 -6.4 0.301
$\blacktriangle$ Gemma-3-12B 12 7.6 59.8 65.9 -7.4 0.316
$\blacktriangle$ Gemma-3-27B 27 7.8 55.7 65.6 -7.1 0.335
$\blacktriangle$ CommandA 111 9.2 60.8 59.2 -8.0 0.254
$\blacktriangle$ NLLB 1 11.4 53.9 55.5 -8.6 0.235
IR-MultiagentMT ? 11.5 53.7 55.5 -8.6 0.238
$\blacktriangle$ TowerPlus-72B[M] 72 13.5 55.0 47.0 -9.9 0.189
$\blacktriangle$ Llama-3.1-8B 8 14.2 50.7 46.1 -9.5 0.176
$\blacktriangle$ ONLINE-G ? 15.8 48.3 48.1 -10.9 0.151
$\blacktriangle$ AyaExpanse-32B 32 17.9 46.2 36.1 -11.7 0.143
$\blacktriangle$ TowerPlus-9B[M] 9 20.3 27.9 9.6 -9.0 0.228
$\blacktriangle$ Qwen2.5-7B ? 7 21.1 36.6 30.8 -12.8 0.122
$\blacktriangle$ CommandR7B 7 22.7 30.6 22.4 -13.8 0.181
$\blacktriangle$ AyaExpanse-8B 8 25.1 27.7 21.5 -16.1 0.16
$\blacktriangle$ EuroLLM-9B[M] 9 27.5 15.5 6.2 -15.2 0.189
$\blacktriangle$ Mistral-7B 7 28.6 19.3 14.4 -18.6 0.175
$\blacktriangle$ EuroLLM-22B-pre.[M] 22 30.0 16.5 12.7 -19.5 0.171
English-German
System Name LP Supported Params. (B) AutoRank $\downarrow$ GEMBA-ESA-CMDA $\uparrow$ GEMBA-ESA-GPT4.1 $\uparrow$ MetricX-24-Hybrid-XL $\uparrow$ XCOMET-XL $\uparrow$
Shy-hunyuan-MT 7 1.0 84.3 90.6 -3.1 0.703
CommandA-WMT 111 2.2 82.8 89.0 -3.0 0.686
$\blacktriangle$ GPT-4.1 ? 3.2 84.6 91.4 -3.6 0.671
$\blacktriangle$ Gemini-2.5-Pro ? 3.4 84.0 91.7 -3.5 0.665
$\blacktriangle$ DeepSeek-V3 ? 671 3.6 84.0 90.0 -3.6 0.671
$\blacktriangle$ Mistral-Medium ? 3.8 83.6 88.2 -3.6 0.676
GemTrans 27 4.7 78.7 84.8 -3.1 0.672
$\blacktriangle$ CommandA 111 4.8 82.3 87.1 -3.8 0.672
$\blacktriangle$ ONLINE-B ? 5.4 77.7 83.4 -3.3 0.678
$\blacktriangle$ Claude-4 ? 5.5 81.4 86.9 -3.9 0.669
UvA-MT 12 5.9 77.5 83.3 -3.6 0.679
$\blacktriangle$ Qwen3-235B 235 6.2 80.0 85.4 -3.7 0.659
$\blacktriangle$ AyaExpanse-32B 32 6.5 78.2 83.8 -3.8 0.669
$\blacktriangle$ Llama-4-Maverick 400 7.0 79.0 83.5 -3.9 0.663
$\blacktriangle$ ONLINE-W ? ? 8.0 76.4 80.9 -3.9 0.664
$\blacktriangle$ TowerPlus-9B[M] 9 8.2 76.0 80.0 -3.9 0.667
$\blacktriangle$ TowerPlus-72B[M] 72 8.4 76.0 80.2 -4.0 0.665
TranssionTranslate ? ? 8.7 73.5 78.1 -3.4 0.653
$\blacktriangle$ EuroLLM-22B-pre.[M] 22 8.7 75.4 79.0 -4.1 0.669
SalamandraTA 8 9.8 72.4 75.4 -3.8 0.663
IRB-MT 12 9.8 74.8 79.0 -3.7 0.63
$\blacktriangle$ Gemma-3-12B 12 12.2 73.0 76.2 -4.4 0.633
$\blacktriangle$ AyaExpanse-8B 8 12.2 70.1 75.3 -4.3 0.644
$\blacktriangle$ EuroLLM-9B[M] 9 12.4 70.5 73.6 -4.5 0.654
IR-MultiagentMT ? 12.9 71.9 77.2 -4.7 0.63
$\blacktriangle$ CommandR7B 7 15.7 67.8 68.8 -4.8 0.628
$\blacktriangle$ Gemma-3-27B 27 17.6 67.4 71.7 -5.1 0.589
$\blacktriangle$ ONLINE-G ? 17.9 66.5 67.7 -5.2 0.609
$\blacktriangle$ Llama-3.1-8B 8 20.9 64.3 62.6 -5.5 0.588
$\blacktriangle$ Qwen2.5-7B 7 23.1 60.0 59.1 -5.5 0.575
$\blacktriangle$ NLLB 1 26.1 58.1 59.3 -6.7 0.573
$\blacktriangle$ Mistral-7B 7 32.0 54.9 50.2 -7.0 0.51
English-Greek
System Name LP Supported Params. (B) AutoRank $\downarrow$ GEMBA-ESA-CMDA $\uparrow$ GEMBA-ESA-GPT4.1 $\uparrow$ MetricX-24-Hybrid-XL $\uparrow$ XCOMET-XL $\uparrow$
Shy-hunyuan-MT 7 1.0 80.3 85.8 -5.3 0.601
$\blacktriangle$ Gemini-2.5-Pro ? 1.9 84.3 88.7 -6.2 0.529
CommandA-WMT 111 2.1 79.9 84.1 -5.7 0.56
$\blacktriangle$ GPT-4.1 ? 2.4 82.6 87.1 -6.4 0.528
GemTrans 27 3.2 74.3 78.9 -5.4 0.543
UvA-MT ? 12 4.0 73.3 77.6 -6.4 0.545
$\blacktriangle$ CommandA 111 4.3 77.4 80.7 -7.1 0.509
$\blacktriangle$ Claude-4 ? ? 4.3 79.0 82.2 -7.3 0.496
SalamandraTA 8 4.7 71.0 75.4 -6.3 0.524
$\blacktriangle$ Mistral-Medium ? ? 5.1 73.5 78.2 -7.0 0.498
$\blacktriangle$ ONLINE-B ? 5.5 71.2 75.0 -6.7 0.495
$\blacktriangle$ ONLINE-W ? ? 5.5 74.0 77.3 -7.4 0.487
$\blacktriangle$ AyaExpanse-32B 32 5.6 72.1 75.8 -7.2 0.494
IRB-MT 12 5.9 70.3 73.9 -6.9 0.486
$\blacktriangle$ DeepSeek-V3 ? 671 6.4 69.9 74.7 -7.6 0.48
$\blacktriangle$ EuroLLM-22B-pre.[M] 22 6.6 69.3 72.1 -7.4 0.482
$\blacktriangle$ Llama-4-Maverick 400 6.6 71.4 74.1 -7.9 0.471
TranssionTranslate ? ? 6.7 67.7 71.5 -6.9 0.468
$\blacktriangle$ Qwen3-235B 235 7.6 67.0 69.8 -7.6 0.455
$\blacktriangle$ AyaExpanse-8B 8 8.0 65.1 67.7 -7.8 0.46
$\blacktriangle$ EuroLLM-9B[M] 9 8.7 62.7 66.1 -8.1 0.454
IR-MultiagentMT ? 9.1 65.6 67.2 -8.6 0.419
$\blacktriangle$ Gemma-3-12B 12 9.9 60.6 62.9 -8.9 0.436
$\blacktriangle$ Gemma-3-27B 27 12.0 54.9 56.9 -9.7 0.411
$\blacktriangle$ ONLINE-G ? 13.2 58.9 60.1 -10.9 0.333
$\blacktriangle$ NLLB 1 13.4 55.1 57.5 -11.1 0.373
$\blacktriangle$ CommandR7B 7 17.6 27.9 17.5 -9.9 0.487
$\blacktriangle$ Llama-3.1-8B 8 19.0 44.8 41.7 -13.2 0.254
$\blacktriangle$ TowerPlus-72B[M] 72 22.4 36.5 33.6 -14.8 0.202
$\blacktriangle$ TowerPlus-9B[M] 9 26.5 26.8 22.9 -16.7 0.148
$\blacktriangle$ Qwen2.5-7B ? 7 29.8 22.1 20.0 -20.0 0.109
$\blacktriangle$ Mistral-7B 7 32.0 19.2 14.3 -22.7 0.135
English-Persian
System Name LP Supported Params. (B) AutoRank $\downarrow$ GEMBA-ESA-CMDA $\uparrow$ GEMBA-ESA-GPT4.1 $\uparrow$ MetricX-24-Hybrid-XL $\uparrow$ XCOMET-XL $\uparrow$
Shy-hunyuan-MT 7 1.0 80.4 84.1 -4.6 0.553
$\blacktriangle$ Gemini-2.5-Pro ? 1.7 82.4 88.4 -5.2 0.476
$\blacktriangle$ GPT-4.1 ? 2.3 81.1 85.4 -5.4 0.47
CommandA-WMT 111 2.6 77.4 80.4 -5.0 0.497
GemTrans 27 2.9 74.1 77.0 -4.5 0.502
$\blacktriangle$ DeepSeek-V3 ? 671 3.3 78.1 80.6 -5.5 0.456
UvA-MT ? 12 3.7 73.2 76.2 -5.3 0.489
$\blacktriangle$ Gemma-3-27B 27 3.8 75.8 79.0 -5.6 0.453
$\blacktriangle$ Mistral-Medium ? ? 3.9 75.5 78.7 -5.6 0.453
$\blacktriangle$ Claude-4 ? ? 4.4 77.5 79.6 -6.3 0.427
$\blacktriangle$ CommandA 111 4.6 74.0 77.1 -6.0 0.439
$\blacktriangle$ ONLINE-B ? 4.8 70.7 72.3 -5.4 0.458
IRB-MT 12 5.1 71.7 73.1 -5.6 0.432
$\blacktriangle$ Llama-4-Maverick 400 5.1 72.3 75.8 -6.0 0.425
TranssionTranslate ? ? 5.6 68.5 69.3 -5.5 0.438
$\blacktriangle$ Gemma-3-12B 12 5.7 71.0 72.5 -6.1 0.417
$\blacktriangle$ AyaExpanse-32B 32 5.7 70.4 72.3 -6.1 0.425
$\blacktriangle$ Qwen3-235B 235 7.8 64.1 66.9 -6.6 0.378
IR-MultiagentMT ? 8.7 63.8 63.8 -7.1 0.359
$\blacktriangle$ AyaExpanse-8B 8 8.8 62.1 62.5 -7.0 0.369
$\blacktriangle$ CommandR7B 7 12.7 55.1 49.5 -8.9 0.312
$\blacktriangle$ ONLINE-G ? 13.4 54.6 53.2 -9.3 0.255
$\blacktriangle$ NLLB 1 13.8 52.5 52.4 -9.6 0.27
$\blacktriangle$ Llama-3.1-8B 8 13.8 51.5 49.2 -8.9 0.261
$\blacktriangle$ TowerPlus-72B[M] 72 16.6 45.6 43.8 -10.3 0.203
$\blacktriangle$ TowerPlus-9B[M] 9 20.2 37.7 32.8 -12.0 0.16
$\blacktriangle$ Qwen2.5-7B ? 7 21.6 32.4 32.0 -12.7 0.134
$\blacktriangle$ EuroLLM-22B-pre.[M] 22 28.3 21.2 16.1 -18.8 0.165
$\blacktriangle$ Mistral-7B 7 28.6 21.9 17.6 -19.0 0.131
$\blacktriangle$ EuroLLM-9B[M] 9 30.0 14.5 9.8 -19.7 0.185
English-Hindi
System Name LP Supported Params. (B) AutoRank $\downarrow$ GEMBA-ESA-CMDA $\uparrow$ GEMBA-ESA-GPT4.1 $\uparrow$ MetricX-24-Hybrid-XL $\uparrow$ XCOMET-XL $\uparrow$
Shy-hunyuan-MT 7 1.0 77.0 82.3 -5.1 0.44
$\blacktriangle$ Gemini-2.5-Pro ? 1.9 78.3 86.3 -5.7 0.376
GemTrans 27 2.6 72.4 78.4 -5.2 0.397
$\blacktriangle$ GPT-4.1 ? 2.7 75.7 84.5 -5.9 0.372
$\blacktriangle$ DeepSeek-V3 ? 671 3.0 76.2 82.4 -5.9 0.36
CommandA-WMT 111 3.2 73.6 79.0 -5.6 0.375
UvA-MT ? 12 4.4 70.8 77.6 -6.0 0.355
$\blacktriangle$ Claude-4 ? 4.8 73.7 78.3 -6.6 0.334
$\blacktriangle$ Gemma-3-27B 27 5.2 71.9 76.6 -6.3 0.319
IRB-MT 12 5.3 69.8 74.3 -6.1 0.33
$\blacktriangle$ ONLINE-B ? 5.6 68.0 74.5 -6.2 0.331
$\blacktriangle$ CommandA 111 5.8 71.0 74.9 -6.6 0.314
TranssionTranslate ? ? 6.5 64.7 70.3 -6.1 0.326
$\blacktriangle$ Llama-4-Maverick 400 6.7 68.5 73.4 -6.7 0.296
$\blacktriangle$ Qwen3-235B 235 6.8 67.8 72.1 -6.6 0.298
$\blacktriangle$ Mistral-Medium ? ? 6.9 67.2 71.7 -6.9 0.322
$\blacktriangle$ Gemma-3-12B 12 7.1 66.8 70.1 -6.8 0.309
$\blacktriangle$ TowerPlus-9B[M] 9 7.5 67.1 70.8 -7.0 0.287
$\blacktriangle$ AyaExpanse-32B 32 8.1 65.6 70.4 -7.1 0.27
$\blacktriangle$ TowerPlus-72B[M] 72 9.3 63.3 66.3 -7.4 0.264
IR-MultiagentMT ? 10.1 61.9 62.9 -7.6 0.251
$\blacktriangle$ EuroLLM-22B-pre.[M] 22 10.7 59.7 61.2 -7.7 0.259
$\blacktriangle$ AyaExpanse-8B 8 10.8 59.3 60.6 -7.7 0.254
$\blacktriangle$ EuroLLM-9B[M] 9 11.6 53.6 54.4 -7.8 0.3
$\blacktriangle$ Llama-3.1-8B 8 13.8 54.8 54.2 -8.6 0.195
$\blacktriangle$ NLLB 1 14.4 55.2 55.2 -9.4 0.199
$\blacktriangle$ ONLINE-G ? 15.4 54.5 51.8 -9.6 0.176
$\blacktriangle$ CommandR7B 7 15.9 49.6 49.6 -9.3 0.18
$\blacktriangle$ Qwen2.5-7B ? 7 24.8 30.0 32.5 -12.8 0.107
$\blacktriangle$ Mistral-7B 7 30.0 25.0 23.2 -16.6 0.126
English-Indonesian
System Name LP Supported Params. (B) AutoRank $\downarrow$ GEMBA-ESA-CMDA $\uparrow$ GEMBA-ESA-GPT4.1 $\uparrow$ MetricX-24-Hybrid-XL $\uparrow$ XCOMET-XL $\uparrow$
Shy-hunyuan-MT 7 1.0 83.2 87.1 -4.4 0.677
$\blacktriangle$ Gemini-2.5-Pro ? 2.8 83.0 89.3 -5.6 0.576
$\blacktriangle$ GPT-4.1 ? 3.6 81.6 87.9 -5.9 0.564
GemTrans 27 3.7 76.4 80.8 -4.7 0.622
CommandA-WMT 111 4.0 78.2 83.7 -5.5 0.592
$\blacktriangle$ DeepSeek-V3 ? 671 4.1 81.2 85.1 -5.9 0.558
$\blacktriangle$ Qwen3-235B 235 4.3 79.8 84.2 -6.0 0.566
UvA-MT ? 12 4.4 78.0 83.2 -5.9 0.584
$\blacktriangle$ Mistral-Medium ? ? 5.1 78.2 83.6 -6.3 0.549
$\blacktriangle$ Gemma-3-27B 27 5.4 78.3 83.1 -6.4 0.531
IRB-MT 12 5.5 75.8 80.6 -5.8 0.548
$\blacktriangle$ Claude-4 ? 5.9 78.8 82.8 -6.9 0.514
$\blacktriangle$ Gemma-3-12B 12 6.6 75.3 81.1 -6.8 0.515
$\blacktriangle$ ONLINE-B ? 7.0 72.7 76.7 -6.3 0.528
$\blacktriangle$ Llama-4-Maverick 400 7.3 74.0 78.5 -6.9 0.507
$\blacktriangle$ CommandA 111 7.5 74.9 77.7 -7.1 0.498
$\blacktriangle$ AyaExpanse-32B 32 7.8 72.7 76.9 -6.9 0.5
$\blacktriangle$ ONLINE-W ? ? 8.2 69.8 73.9 -6.7 0.522
TranssionTranslate ? ? 8.5 68.7 72.7 -6.3 0.498
$\blacktriangle$ TowerPlus-72B[M] 72 9.1 70.2 74.6 -7.5 0.479
$\blacktriangle$ AyaExpanse-8B 8 9.3 68.6 72.1 -7.2 0.487
IR-MultiagentMT ? 9.8 68.4 72.6 -7.5 0.464
$\blacktriangle$ ONLINE-G ? 12.9 63.4 65.9 -8.7 0.409
$\blacktriangle$ Llama-3.1-8B 8 13.6 62.2 62.4 -9.0 0.417
$\blacktriangle$ Qwen2.5-7B ? 7 13.8 60.2 61.6 -8.6 0.412
$\blacktriangle$ CommandR7B 7 16.6 57.6 53.0 -10.2 0.392
$\blacktriangle$ NLLB 1 17.3 57.3 57.7 -10.9 0.333
$\blacktriangle$ TowerPlus-9B[M] 9 18.7 52.0 50.2 -10.6 0.339
$\blacktriangle$ EuroLLM-22B-pre.[M] 22 25.5 40.6 39.4 -13.7 0.214
$\blacktriangle$ Mistral-7B 7 25.6 43.1 40.1 -14.2 0.197
$\blacktriangle$ EuroLLM-9B[M] 9 31.0 26.4 20.2 -16.0 0.275
English-Kannada
System Name LP Supported Params. (B) AutoRank $\downarrow$ GEMBA-ESA-CMDA $\uparrow$ GEMBA-ESA-GPT4.1 $\uparrow$ MetricX-24-Hybrid-XL $\uparrow$ XCOMET-XL $\uparrow$
Shy-hunyuan-MT 7 1.0 64.0 78.8 -6.0 0.446
$\blacktriangle$ Gemini-2.5-Pro ? 2.2 62.5 81.6 -6.3 0.399
$\blacktriangle$ Claude-4 ? ? 5.0 61.3 76.1 -7.6 0.333
GemTrans 27 5.2 58.7 67.3 -6.7 0.358
$\blacktriangle$ GPT-4.1 ? 5.9 60.2 71.3 -7.9 0.327
$\blacktriangle$ Mistral-Medium ? ? 6.5 59.8 69.2 -8.0 0.312
$\blacktriangle$ Qwen3-235B 235 6.7 60.1 67.4 -7.9 0.305
$\blacktriangle$ DeepSeek-V3 ? 671 6.7 57.3 69.9 -8.3 0.325
CommandA-WMT 111 7.5 59.7 64.5 -8.4 0.295
$\blacktriangle$ ONLINE-B ? 7.7 57.0 66.0 -7.9 0.289
$\blacktriangle$ Gemma-3-27B 27 7.8 57.2 66.1 -8.3 0.294
TranssionTranslate ? ? 8.1 56.1 63.3 -7.8 0.286
$\blacktriangle$ Llama-4-Maverick 400 8.1 58.1 66.5 -8.4 0.27
UvA-MT ? 12 8.8 53.3 60.5 -8.6 0.308
IRB-MT 12 11.0 52.4 57.6 -9.3 0.239
$\blacktriangle$ NLLB 1 12.1 52.3 54.2 -9.8 0.215
$\blacktriangle$ ONLINE-G ? 13.3 52.5 51.9 -10.5 0.186
$\blacktriangle$ Gemma-3-12B 12 13.4 46.1 49.4 -10.4 0.244
$\blacktriangle$ CommandA 111 14.2 54.3 48.1 -11.6 0.175
$\blacktriangle$ TowerPlus-9B[M] 9 18.1 32.8 2.7 -8.8 0.281
$\blacktriangle$ Llama-3.1-8B 8 19.1 44.1 35.8 -13.4 0.12
IR-MultiagentMT ? 19.5 40.1 36.4 -13.9 0.149
$\blacktriangle$ AyaExpanse-32B 32 23.9 34.3 25.5 -18.6 0.157
$\blacktriangle$ CommandR7B 7 25.6 17.1 10.6 -16.4 0.222
$\blacktriangle$ TowerPlus-72B[M] 72 25.8 22.4 16.5 -18.3 0.197
$\blacktriangle$ AyaExpanse-8B 8 28.1 19.8 14.3 -20.5 0.174
$\blacktriangle$ EuroLLM-9B[M] 9 28.2 13.1 2.9 -16.9 0.179
$\blacktriangle$ EuroLLM-22B-pre.[M] 22 28.7 12.8 5.0 -18.3 0.184
$\blacktriangle$ Mistral-7B 7 29.9 7.4 4.6 -19.2 0.199
$\blacktriangle$ Qwen2.5-7B ? 7 30.0 15.1 10.2 -21.4 0.163
English-Lithuanian
System Name LP Supported Params. (B) AutoRank $\downarrow$ GEMBA-ESA-CMDA $\uparrow$ GEMBA-ESA-GPT4.1 $\uparrow$ MetricX-24-Hybrid-XL $\uparrow$ XCOMET-XL $\uparrow$
Shy-hunyuan-MT 7 1.0 77.6 84.1 -6.3 0.569
$\blacktriangle$ Gemini-2.5-Pro ? 2.3 76.1 87.3 -7.2 0.502
$\blacktriangle$ GPT-4.1 ? 2.9 75.3 84.8 -7.6 0.5
CommandA-WMT 111 4.5 72.6 72.4 -7.8 0.506
GemTrans 27 4.5 70.2 71.7 -6.8 0.505
$\blacktriangle$ ONLINE-B ? 5.4 69.1 70.9 -7.7 0.487
$\blacktriangle$ Claude-4 ? ? 5.8 71.8 75.6 -9.3 0.455
SalamandraTA 8 6.1 66.9 67.0 -7.9 0.496
$\blacktriangle$ ONLINE-W ? ? 6.7 67.6 69.5 -9.1 0.467
TranssionTranslate ? ? 6.7 66.3 67.6 -7.9 0.454
$\blacktriangle$ Gemma-3-27B 27 6.9 69.8 68.8 -9.0 0.434
$\blacktriangle$ Llama-4-Maverick 400 6.9 69.1 71.5 -9.3 0.43
UvA-MT ? 12 7.1 68.5 63.7 -9.0 0.472
$\blacktriangle$ Qwen3-235B 235 7.8 67.8 66.1 -9.1 0.414
$\blacktriangle$ EuroLLM-22B-pre.[M] 22 8.3 64.7 66.1 -9.7 0.434
IRB-MT 12 8.9 66.0 61.2 -9.5 0.402
$\blacktriangle$ EuroLLM-9B[M] 9 9.3 61.0 57.5 -9.5 0.455
$\blacktriangle$ DeepSeek-V3 ? 671 9.7 60.4 60.8 -9.8 0.418
$\blacktriangle$ Gemma-3-12B 12 10.2 66.0 58.3 -10.7 0.368
IR-MultiagentMT ? 10.5 63.2 59.8 -10.9 0.374
$\blacktriangle$ Mistral-Medium ? ? 10.8 64.5 56.7 -10.9 0.362
$\blacktriangle$ CommandA 111 11.0 65.1 55.7 -11.4 0.361
$\blacktriangle$ ONLINE-G ? 15.5 55.7 52.4 -14.2 0.259
$\blacktriangle$ NLLB 1 15.9 53.0 49.3 -14.2 0.283
$\blacktriangle$ AyaExpanse-32B 32 22.3 46.0 32.7 -17.9 0.143
$\blacktriangle$ TowerPlus-72B[M] 72 23.6 41.4 28.6 -18.1 0.141
$\blacktriangle$ Llama-3.1-8B 8 23.9 40.4 31.2 -18.6 0.125
$\blacktriangle$ TowerPlus-9B[M] 9 26.3 34.3 18.3 -19.8 0.157
$\blacktriangle$ CommandR7B 7 27.0 21.5 4.6 -17.7 0.274
$\blacktriangle$ AyaExpanse-8B 8 29.3 25.8 17.5 -22.7 0.141
$\blacktriangle$ Qwen2.5-7B ? 7 29.7 24.1 19.0 -22.7 0.116
$\blacktriangle$ Mistral-7B 7 32.0 16.0 11.4 -24.0 0.14
English-Marathi
System Name LP Supported Params. (B) AutoRank $\downarrow$ GEMBA-ESA-CMDA $\uparrow$ GEMBA-ESA-GPT4.1 $\uparrow$ MetricX-24-Hybrid-XL $\uparrow$ XCOMET-XL $\uparrow$
Shy-hunyuan-MT 7 1.0 70.8 81.6 -5.8 0.248
$\blacktriangle$ Gemini-2.5-Pro ? 2.7 68.1 84.7 -6.2 0.222
GemTrans 27 4.0 67.3 65.2 -5.8 0.224
$\blacktriangle$ GPT-4.1 ? 4.6 67.6 79.4 -6.7 0.196
UvA-MT ? 12 5.2 67.4 72.4 -6.5 0.192
$\blacktriangle$ Claude-4 ? ? 5.5 67.2 76.2 -7.2 0.193
$\blacktriangle$ DeepSeek-V3 ? 671 5.5 67.5 74.7 -7.0 0.19
$\blacktriangle$ Gemma-3-27B 27 5.8 67.1 71.8 -7.0 0.191
$\blacktriangle$ Mistral-Medium ? ? 6.9 66.8 70.6 -7.3 0.171
CommandA-WMT 111 7.2 66.4 64.7 -7.2 0.178
IRB-MT 12 7.3 64.4 68.1 -7.0 0.175
$\blacktriangle$ Llama-4-Maverick 400 7.8 64.0 69.4 -7.5 0.169
$\blacktriangle$ ONLINE-B ? 8.0 62.7 66.0 -7.2 0.172
TranssionTranslate ? ? 8.1 62.6 65.0 -7.1 0.17
$\blacktriangle$ Qwen3-235B 235 8.3 64.1 64.9 -7.5 0.167
$\blacktriangle$ TowerPlus-9B[M] 9 8.9 63.0 7.7 -7.7 0.277
$\blacktriangle$ NLLB 1 12.0 58.3 55.6 -8.7 0.148
IR-MultiagentMT ? 12.1 57.7 55.5 -9.1 0.156
$\blacktriangle$ Gemma-3-12B 12 12.4 52.5 51.8 -9.4 0.189
$\blacktriangle$ ONLINE-G ? 13.2 57.3 54.2 -9.4 0.138
$\blacktriangle$ CommandA 111 13.3 60.9 49.6 -9.7 0.131
$\blacktriangle$ EuroLLM-9B[M] 9 14.1 52.5 10.9 -9.2 0.225
$\blacktriangle$ Llama-3.1-8B 8 17.2 50.4 41.6 -11.3 0.139
$\blacktriangle$ TowerPlus-72B[M] 72 17.8 49.8 30.5 -12.5 0.175
$\blacktriangle$ EuroLLM-22B-pre.[M] 22 18.1 47.0 15.3 -11.8 0.199
$\blacktriangle$ AyaExpanse-32B 32 18.4 49.6 34.7 -13.0 0.163
$\blacktriangle$ AyaExpanse-8B 8 20.8 41.8 27.0 -14.5 0.189
$\blacktriangle$ CommandR7B 7 21.4 36.6 20.0 -13.1 0.187
$\blacktriangle$ Qwen2.5-7B ? 7 27.8 27.6 19.8 -18.0 0.175
$\blacktriangle$ Mistral-7B 7 30.0 25.0 12.5 -17.9 0.146
English-Romanian
System Name LP Supported Params. (B) AutoRank $\downarrow$ GEMBA-ESA-CMDA $\uparrow$ GEMBA-ESA-GPT4.1 $\uparrow$ MetricX-24-Hybrid-XL $\uparrow$ XCOMET-XL $\uparrow$
Shy-hunyuan-MT 7 1.0 83.2 86.3 -5.7 0.651
CommandA-WMT 111 1.8 82.5 86.0 -6.0 0.634
$\blacktriangle$ Gemini-2.5-Pro ? 2.3 85.0 89.3 -6.6 0.586
$\blacktriangle$ GPT-4.1 ? 2.7 83.5 88.2 -6.8 0.597
GemTrans 27 3.2 77.7 80.7 -5.7 0.619
$\blacktriangle$ DeepSeek-V3 ? 671 4.2 80.1 84.4 -6.8 0.574
UvA-MT ? 12 4.7 77.6 80.5 -6.9 0.598
$\blacktriangle$ Mistral-Medium ? ? 5.1 77.7 83.2 -7.2 0.568
$\blacktriangle$ CommandA 111 5.2 79.4 82.9 -7.4 0.563
$\blacktriangle$ Gemma-3-27B 27 5.3 78.9 82.3 -7.4 0.562
$\blacktriangle$ TowerPlus-9B[M] 9 6.0 74.8 79.9 -7.1 0.566
$\blacktriangle$ Claude-4 ? ? 6.1 79.4 82.4 -7.8 0.536
$\blacktriangle$ Qwen3-235B 235 6.2 75.7 78.9 -7.2 0.558
$\blacktriangle$ AyaExpanse-32B 32 6.4 76.4 79.9 -7.5 0.546
IRB-MT 12 6.4 75.4 77.4 -7.0 0.548
SalamandraTA 8 6.5 70.9 75.2 -6.7 0.589
$\blacktriangle$ Llama-4-Maverick 400 6.5 76.4 81.1 -7.6 0.541
$\blacktriangle$ ONLINE-B ? 6.7 73.5 76.9 -7.1 0.556
$\blacktriangle$ Gemma-3-12B 12 7.9 74.3 77.9 -8.0 0.524
$\blacktriangle$ EuroLLM-22B-pre.[M] 22 8.2 71.6 76.4 -7.8 0.533
TranssionTranslate ? ? 9.0 68.4 72.1 -7.3 0.521
$\blacktriangle$ TowerPlus-72B[M] 72 9.3 70.3 73.2 -8.0 0.512
$\blacktriangle$ ONLINE-W ? ? 9.4 72.2 75.4 -8.7 0.51
$\blacktriangle$ AyaExpanse-8B 8 9.8 68.3 71.7 -8.0 0.516
$\blacktriangle$ EuroLLM-9B[M] 9 10.2 68.6 70.6 -8.3 0.512
IR-MultiagentMT ? 16.5 57.6 59.3 -10.0 0.425
$\blacktriangle$ CommandR7B 7 16.9 59.9 54.3 -10.2 0.434
$\blacktriangle$ Llama-3.1-8B 8 18.0 56.9 56.7 -10.3 0.38
$\blacktriangle$ ONLINE-G ? 18.7 59.5 60.6 -11.6 0.359
$\blacktriangle$ NLLB 1 19.3 55.0 57.2 -11.6 0.39
$\blacktriangle$ Mistral-7B 7 28.1 44.6 41.7 -14.0 0.224
$\blacktriangle$ Qwen2.5-7B ? 7 32.0 37.7 34.9 -15.2 0.177
English-Thai
System Name LP Supported Params. (B) AutoRank $\downarrow$ GEMBA-ESA-CMDA $\uparrow$ GEMBA-ESA-GPT4.1 $\uparrow$ MetricX-24-Hybrid-XL $\uparrow$ XCOMET-XL $\uparrow$
Shy-hunyuan-MT 7 1.0 71.3 87.9 -5.1 0.603
$\blacktriangle$ Gemini-2.5-Pro ? 2.2 69.0 90.6 -5.6 0.533
GemTrans 27 2.7 67.9 80.4 -5.4 0.558
UvA-MT ? 12 3.2 69.5 79.7 -6.0 0.54
$\blacktriangle$ GPT-4.1 ? 3.2 69.9 87.2 -6.2 0.489
$\blacktriangle$ Qwen3-235B 235 3.5 68.8 80.9 -6.1 0.51
$\blacktriangle$ DeepSeek-V3 ? 671 3.6 69.6 82.9 -6.3 0.493
$\blacktriangle$ Gemma-3-27B 27 4.1 68.3 82.2 -6.5 0.482
$\blacktriangle$ Mistral-Medium ? ? 4.2 68.9 79.8 -6.6 0.486
$\blacktriangle$ Claude-4 ? ? 4.5 68.5 80.7 -6.8 0.466
IRB-MT 12 4.8 66.2 77.1 -6.4 0.475
$\blacktriangle$ Llama-4-Maverick 400 5.0 67.0 76.2 -6.6 0.463
$\blacktriangle$ ONLINE-B ? 5.0 65.1 72.2 -6.1 0.484
CommandA-WMT 111 5.8 66.8 70.0 -6.8 0.449
TranssionTranslate ? ? 6.3 62.2 67.7 -6.4 0.453
$\blacktriangle$ TowerPlus-72B[M] 72 6.7 64.1 70.3 -7.2 0.424
$\blacktriangle$ Gemma-3-12B 12 9.1 55.1 62.6 -8.1 0.427
IR-MultiagentMT ? 9.5 56.6 56.5 -7.9 0.404
$\blacktriangle$ CommandA 111 11.0 60.6 54.6 -9.4 0.311
$\blacktriangle$ Qwen2.5-7B 7 11.7 56.4 51.1 -9.2 0.319
$\blacktriangle$ Llama-3.1-8B 8 12.6 55.6 51.2 -9.7 0.277
$\blacktriangle$ NLLB 1 14.8 51.2 48.4 -11.3 0.247
$\blacktriangle$ TowerPlus-9B[M] 9 17.5 42.1 34.2 -11.3 0.221
$\blacktriangle$ ONLINE-G ? 21.1 35.8 36.1 -14.9 0.176
$\blacktriangle$ AyaExpanse-32B 32 21.6 36.6 30.5 -14.9 0.154
$\blacktriangle$ Mistral-7B 7 24.6 31.0 24.9 -17.2 0.132
$\blacktriangle$ CommandR7B 7 25.8 27.0 21.6 -18.2 0.156
$\blacktriangle$ AyaExpanse-8B 8 26.4 24.6 20.1 -18.5 0.167
$\blacktriangle$ EuroLLM-22B-pre.[M] 22 29.1 19.0 15.2 -20.9 0.169
$\blacktriangle$ EuroLLM-9B[M] 9 30.0 12.8 7.4 -20.2 0.185
English-Serbian (Latin)
System Name LP Supported Params. (B) AutoRank $\downarrow$ GEMBA-ESA-CMDA $\uparrow$ GEMBA-ESA-GPT4.1 $\uparrow$ MetricX-24-Hybrid-XL $\uparrow$ XCOMET-XL $\uparrow$
Shy-hunyuan-MT 7 1.0 80.1 84.2 -3.4 0.583
Wenyiil 14 2.5 77.8 84.6 -3.8 0.513
Algharb 14 2.8 77.9 86.5 -4.0 0.493
GemTrans 27 2.9 74.6 75.3 -3.4 0.528
$\blacktriangle$ GPT-4.1 ? 2.9 78.6 85.3 -4.1 0.501
$\blacktriangle$ DeepSeek-V3 ? 671 2.9 78.5 80.3 -3.9 0.514
UvA-MT 12 2.9 75.0 75.0 -3.7 0.562
$\blacktriangle$ Gemini-2.5-Pro ? 2.9 77.7 86.9 -4.1 0.488
Yolu 14 3.0 73.0 73.1 -3.4 0.553
$\blacktriangle$ Claude-4 ? ? 4.8 74.5 76.6 -4.5 0.471
SalamandraTA 8 5.0 68.8 68.3 -3.8 0.491
$\blacktriangle$ Llama-4-Maverick 400 6.1 71.3 70.4 -4.7 0.448
IRB-MT 12 6.3 69.0 66.7 -4.3 0.441
$\blacktriangle$ Qwen3-235B 235 6.4 68.8 65.4 -4.3 0.439
IR-MultiagentMT ? 6.8 70.0 66.2 -4.7 0.437
CommandA-WMT 111 7.2 70.4 62.5 -5.6 0.506
$\blacktriangle$ ONLINE-B ? 7.2 69.6 64.6 -5.6 0.497
$\blacktriangle$ Gemma-3-12B 12 7.6 67.8 63.6 -5.0 0.427
$\blacktriangle$ CommandA 111 7.6 67.7 59.9 -4.6 0.41
$\blacktriangle$ Gemma-3-27B 27 9.0 64.1 63.6 -5.8 0.438
$\blacktriangle$ EuroLLM-22B-pre.[M] 22 9.0 60.6 54.4 -4.9 0.43
CUNI-SFT 9 9.4 61.5 53.2 -4.9 0.392
$\blacktriangle$ EuroLLM-9B[M] 9 10.2 58.8 47.7 -5.2 0.42
TranssionTranslate ? ? 10.7 60.8 59.6 -5.9 0.349
TranssionMT 1 11.0 57.3 52.8 -5.4 0.358
$\blacktriangle$ ONLINE-G ? 12.5 57.8 52.9 -6.9 0.375
$\blacktriangle$ TowerPlus-72B[M] 72 12.6 55.7 43.1 -5.5 0.306
$\blacktriangle$ Llama-3.1-8B 8 13.4 54.7 43.8 -6.0 0.29
$\blacktriangle$ AyaExpanse-32B 32 13.8 52.9 40.4 -5.7 0.259
$\blacktriangle$ TowerPlus-9B[M] 9 17.6 43.0 29.2 -6.4 0.181
$\blacktriangle$ Mistral-7B 7 17.6 49.4 37.0 -7.8 0.213
$\blacktriangle$ Qwen2.5-7B ? 7 20.5 39.3 29.0 -8.1 0.144
$\blacktriangle$ AyaExpanse-8B 8 20.7 37.5 25.9 -7.9 0.143
$\blacktriangle$ CommandR7B 7 21.1 38.5 25.7 -8.9 0.203
$\blacktriangle$ NLLB 1 35.0 0.8 0.1 -15.2 0.195
English-Swedish
System Name LP Supported Params. (B) AutoRank $\downarrow$ GEMBA-ESA-CMDA $\uparrow$ GEMBA-ESA-GPT4.1 $\uparrow$ MetricX-24-Hybrid-XL $\uparrow$ XCOMET-XL $\uparrow$
Shy-hunyuan-MT 7 1.0 84.2 91.0 -4.7 0.685
$\blacktriangle$ Gemini-2.5-Pro ? 2.5 83.1 92.3 -5.4 0.638
GemTrans 27 2.9 79.2 85.1 -4.7 0.656
$\blacktriangle$ GPT-4.1 ? 3.2 81.5 91.7 -5.9 0.635
$\blacktriangle$ DeepSeek-V3 ? 671 4.1 81.0 86.8 -5.9 0.621
CommandA-WMT 111 4.4 78.2 81.9 -5.3 0.63
UvA-MT ? 12 4.5 79.0 82.9 -5.7 0.636
$\blacktriangle$ Mistral-Medium ? ? 4.6 80.8 85.6 -6.1 0.614
$\blacktriangle$ Gemma-3-27B 27 5.0 79.5 84.7 -6.1 0.61
$\blacktriangle$ Claude-4 ? ? 5.3 80.9 85.4 -6.6 0.601
IRB-MT 12 5.8 76.3 80.4 -5.8 0.606
$\blacktriangle$ TowerPlus-9B[M] 9 6.0 77.0 81.3 -6.2 0.602
$\blacktriangle$ ONLINE-B ? 6.1 76.2 80.5 -6.1 0.599
SalamandraTA 8 6.1 75.0 78.0 -6.0 0.621
$\blacktriangle$ Llama-4-Maverick 400 6.2 78.4 81.8 -6.6 0.591
$\blacktriangle$ ONLINE-W ? ? 7.2 75.6 80.7 -7.0 0.591
$\blacktriangle$ TowerPlus-72B[M] 72 7.5 75.2 77.6 -6.8 0.58
$\blacktriangle$ Qwen3-235B 235 8.2 73.9 75.3 -6.8 0.571
$\blacktriangle$ EuroLLM-22B-pre.[M] 22 8.3 74.4 76.6 -7.2 0.568
IR-MultiagentMT ? 8.3 74.6 76.9 -7.2 0.564
TranssionTranslate ? ? 8.4 69.7 75.7 -6.2 0.563
$\blacktriangle$ CommandA 111 8.9 74.6 75.0 -7.3 0.551
$\blacktriangle$ EuroLLM-9B[M] 9 9.8 70.8 72.3 -7.3 0.555
$\blacktriangle$ Gemma-3-12B 12 11.4 67.2 69.2 -7.7 0.528
$\blacktriangle$ Llama-3.1-8B 8 14.6 63.8 61.1 -8.8 0.483
$\blacktriangle$ ONLINE-G ? 17.7 61.8 59.9 -10.6 0.422
$\blacktriangle$ NLLB 1 18.1 58.9 58.2 -10.5 0.436
$\blacktriangle$ Mistral-7B 7 21.0 56.2 50.3 -11.2 0.374
$\blacktriangle$ AyaExpanse-32B 32 21.1 55.1 49.6 -11.1 0.376
$\blacktriangle$ Qwen2.5-7B ? 7 26.0 47.5 42.0 -12.9 0.304
$\blacktriangle$ CommandR7B 7 27.8 41.0 31.7 -12.7 0.316
$\blacktriangle$ AyaExpanse-8B 8 32.0 40.1 33.7 -15.5 0.211
English-Turkish
System Name LP Supported Params. (B) AutoRank $\downarrow$ GEMBA-ESA-CMDA $\uparrow$ GEMBA-ESA-GPT4.1 $\uparrow$ MetricX-24-Hybrid-XL $\uparrow$ XCOMET-XL $\uparrow$
Shy-hunyuan-MT 7 1.0 81.4 85.2 -7.2 0.542
$\blacktriangle$ Gemini-2.5-Pro ? 2.7 82.7 87.9 -8.4 0.462
$\blacktriangle$ GPT-4.1 ? 3.0 83.1 86.1 -8.6 0.465
CommandA-WMT 111 3.3 77.9 80.2 -7.8 0.491
GemTrans 27 3.3 74.6 78.8 -7.4 0.506
$\blacktriangle$ DeepSeek-V3 ? 671 3.4 81.2 84.5 -8.6 0.461
$\blacktriangle$ Mistral-Medium ? ? 5.3 76.0 78.7 -9.0 0.44
$\blacktriangle$ Claude-4 ? 5.5 77.7 80.1 -9.4 0.424
UvA-MT ? 12 5.6 72.6 76.0 -8.8 0.46
$\blacktriangle$ ONLINE-W ? ? 6.2 73.7 76.1 -9.1 0.431
$\blacktriangle$ ONLINE-B ? 7.1 72.4 72.5 -9.3 0.414
IRB-MT 12 7.2 69.9 71.8 -8.9 0.415
$\blacktriangle$ Llama-4-Maverick 400 7.3 72.5 74.7 -9.8 0.409
TranssionTranslate ? ? 7.6 68.1 71.6 -9.1 0.413
$\blacktriangle$ Qwen3-235B 235 7.7 69.4 71.2 -9.3 0.408
$\blacktriangle$ CommandA 111 8.1 71.5 72.2 -9.9 0.393
$\blacktriangle$ Gemma-3-12B 12 8.7 68.5 69.4 -9.8 0.391
$\blacktriangle$ EuroLLM-22B-pre.[M] 22 9.0 66.1 69.1 -9.9 0.397
$\blacktriangle$ Gemma-3-27B 27 9.1 66.9 69.6 -10.1 0.394
$\blacktriangle$ AyaExpanse-32B 32 9.9 64.5 66.1 -10.2 0.383
$\blacktriangle$ EuroLLM-9B[M] 9 10.8 59.6 60.5 -10.2 0.409
IR-MultiagentMT ? 10.9 64.1 65.3 -10.6 0.351
$\blacktriangle$ AyaExpanse-8B 8 13.0 58.6 58.7 -11.0 0.325
$\blacktriangle$ TowerPlus-72B[M] 72 13.5 58.5 56.7 -11.3 0.325
$\blacktriangle$ ONLINE-G ? 14.3 58.0 58.6 -11.9 0.294
$\blacktriangle$ NLLB 1 15.5 53.3 55.3 -12.4 0.304
$\blacktriangle$ Llama-3.1-8B 8 17.8 51.1 48.4 -12.9 0.248
$\blacktriangle$ CommandR7B 7 18.0 48.4 42.9 -12.8 0.291
$\blacktriangle$ TowerPlus-9B[M] 9 22.1 43.6 36.9 -14.6 0.192
$\blacktriangle$ Qwen2.5-7B ? 7 22.7 41.2 38.5 -14.9 0.174
$\blacktriangle$ Mistral-7B 7 31.0 27.1 22.2 -20.2 0.138
English-Vietnamese
System Name LP Supported Params. (B) AutoRank $\downarrow$ GEMBA-ESA-CMDA $\uparrow$ GEMBA-ESA-GPT4.1 $\uparrow$ MetricX-24-Hybrid-XL $\uparrow$ XCOMET-XL $\uparrow$
Shy-hunyuan-MT 7 1.0 83.1 87.3 -4.5 0.623
$\blacktriangle$ Gemini-2.5-Pro ? 2.7 82.3 88.6 -5.6 0.539
CommandA-WMT 111 2.7 78.4 83.2 -4.9 0.577
$\blacktriangle$ GPT-4.1 ? 2.8 82.9 88.1 -5.7 0.533
$\blacktriangle$ DeepSeek-V3 ? 671 3.2 81.5 85.5 -5.7 0.533
$\blacktriangle$ Qwen3-235B 235 3.3 79.9 84.1 -5.5 0.539
GemTrans 27 3.4 74.8 80.3 -4.8 0.572
UvA-MT ? 12 3.7 77.0 80.8 -5.5 0.559
$\blacktriangle$ Mistral-Medium ? ? 3.7 78.7 83.8 -5.8 0.53
$\blacktriangle$ Claude-4 ? ? 5.0 78.2 81.5 -6.7 0.494
IRB-MT 12 5.1 74.1 77.7 -5.8 0.506
$\blacktriangle$ AyaExpanse-32B 32 5.8 72.3 76.9 -6.3 0.498
$\blacktriangle$ Llama-4-Maverick 400 6.6 72.2 76.4 -6.9 0.47
$\blacktriangle$ ONLINE-B ? 6.6 70.9 74.4 -6.6 0.478
TranssionTranslate ? ? 7.2 66.4 70.7 -6.1 0.476
$\blacktriangle$ AyaExpanse-8B 8 7.8 66.3 70.1 -6.7 0.465
$\blacktriangle$ Gemma-3-12B 12 8.1 67.2 70.9 -7.2 0.448
$\blacktriangle$ CommandA 111 8.7 66.8 69.2 -7.7 0.442
IR-MultiagentMT ? 8.7 67.0 70.0 -7.5 0.424
$\blacktriangle$ TowerPlus-72B[M] 72 8.8 65.3 65.4 -7.3 0.46
$\blacktriangle$ Gemma-3-27B 27 9.9 62.8 65.5 -7.9 0.42
$\blacktriangle$ Qwen2.5-7B 7 10.8 61.4 61.2 -8.3 0.41
$\blacktriangle$ Llama-3.1-8B 8 11.8 59.3 60.7 -8.9 0.385
$\blacktriangle$ CommandR7B 7 13.1 55.7 52.0 -9.6 0.406
$\blacktriangle$ NLLB 1 15.5 54.0 53.9 -11.4 0.303
$\blacktriangle$ TowerPlus-9B[M] 9 16.8 46.2 42.1 -10.7 0.319
$\blacktriangle$ ONLINE-G ? 17.4 52.5 51.0 -12.6 0.238
$\blacktriangle$ Mistral-7B 7 24.4 33.7 33.6 -15.9 0.139
$\blacktriangle$ EuroLLM-9B[M] 9 27.3 18.8 9.4 -17.9 0.327
$\blacktriangle$ EuroLLM-22B-pre.[M] 22 30.0 22.2 20.8 -20.5 0.113
Appendix A Metrics correlations
To examine how the metrics used for AutoRankcorrelate with each other, we calculated the Pearson correlation between paragraph-level scores for all systems, resulting in a sample size of around 14k scores per each language pair.
The results show that GEMBA-ESA on CmdA and GPT-4.1 exhibit the highest correlations for almost all languages. In contrast, the weakest correlations are generally observed between xComet and both GEMBA-ESA variants.
When examining results by language pair, Bhojpuri, Maasai, and Marathi show the lowest correlations. This is why we use chrF++ for the first two language pairs. Unfortunately, no reference translations are available for Marathi, so we must rely on QE metrics for its evaluation.
Kiwi G-CmdA Kiwi G-GPT Kiwi MetX Kiwi xComet G-CmdA G-GPT G-CmdA MetX G-CmdA xComet G-GPT MetX G-GPT xComet MetX xComet
cs-de_DE 0.441 0.484 0.541 0.709 0.732 0.583 0.403 0.636 0.436 0.560
cs-uk_UA 0.531 0.600 0.696 0.794 0.708 0.571 0.517 0.654 0.573 0.710
en-ar_EG 0.610 0.573 0.750 0.494 0.740 0.624 0.350 0.605 0.268 0.565
en-bho_IN 0.465 0.093 0.517 0.030 0.503 0.621 0.051 0.428 -0.008 0.194
en-bn_BD 0.742 0.752 0.822 0.498 0.802 0.735 0.435 0.730 0.448 0.584
en-cs_CZ 0.617 0.696 0.728 0.747 0.757 0.642 0.533 0.682 0.535 0.712
en-de_DE 0.481 0.546 0.612 0.789 0.742 0.578 0.350 0.593 0.358 0.559
en-el_GR 0.736 0.777 0.787 0.691 0.863 0.716 0.542 0.743 0.544 0.741
en-et_EE 0.783 0.837 0.825 0.720 0.787 0.736 0.583 0.795 0.655 0.802
en-fa_IR 0.814 0.834 0.862 0.703 0.852 0.785 0.596 0.793 0.589 0.689
en-hi_IN 0.651 0.663 0.654 0.443 0.754 0.658 0.432 0.681 0.459 0.634
en-id_ID 0.696 0.777 0.705 0.680 0.775 0.633 0.542 0.653 0.552 0.775
en-is_IS 0.787 0.811 0.839 0.659 0.756 0.713 0.495 0.787 0.620 0.741
en-it_IT 0.549 0.596 0.691 0.780 0.735 0.566 0.470 0.583 0.456 0.716
en-ja_JP 0.644 0.668 0.717 0.691 0.752 0.626 0.543 0.637 0.496 0.715
en-kn_IN 0.796 0.778 0.826 0.379 0.790 0.714 0.324 0.703 0.375 0.563
en-ko_KR 0.645 0.667 0.699 0.680 0.774 0.643 0.580 0.648 0.547 0.738
en-lt_LT 0.798 0.837 0.858 0.726 0.828 0.755 0.556 0.783 0.601 0.762
en-mas_KE 0.694 0.325 0.403 0.124 0.460 0.406 0.223 0.096 -0.085 0.533
en-mr_IN 0.738 0.622 0.785 0.179 0.610 0.685 0.124 0.595 0.034 0.320
en-ro_RO 0.634 0.707 0.748 0.796 0.753 0.619 0.546 0.648 0.561 0.762
en-ru_RU 0.580 0.647 0.677 0.731 0.707 0.534 0.499 0.575 0.500 0.742
en-sr_Cyrl_RS 0.699 0.775 0.714 0.743 0.737 0.577 0.577 0.655 0.664 0.696
en-sr_Latn_RS 0.731 0.789 0.724 0.691 0.797 0.672 0.532 0.661 0.564 0.610
en-sv_SE 0.662 0.738 0.777 0.830 0.780 0.634 0.573 0.706 0.641 0.798
en-th_TH 0.821 0.845 0.837 0.667 0.831 0.775 0.585 0.797 0.639 0.735
en-tr_TR 0.704 0.758 0.713 0.649 0.782 0.619 0.498 0.642 0.516 0.738
en-uk_UA 0.646 0.704 0.745 0.763 0.752 0.594 0.550 0.643 0.568 0.771
en-vi_VN 0.714 0.762 0.762 0.641 0.827 0.685 0.507 0.698 0.522 0.743
en-zh_CN 0.557 0.633 0.653 0.653 0.688 0.584 0.525 0.584 0.518 0.744
ja-zh_CN 0.508 0.553 0.658 0.735 0.779 0.639 0.532 0.639 0.545 0.718